### v1

# MixGRPO: Unlocking Flow-based GRPO Efficiency with Mixed ODE-SDE

Abstract
Although GRPO substantially enhances flow matching models in human preference alignment of image generation, methods such as FlowGRPO still exhibit inefficiency due to the necessity of sampling and optimizing over all denoising steps specified by the Markov Decision Process (MDP). In this paper, we propose MixGRPO, a novel framework that leverages the flexibility of mixed sampling strategies through the integration of stochastic differential equations (SDE) and ordinary differential equations (ODE). This streamlines the optimization process within the MDP to improve efficiency and boost performance. Specifically, MixGRPO introduces a sliding window mechanism, using SDE sampling and GRPO-guided optimization only within the window, while applying ODE sampling outside. This design confines sampling randomness to the time-steps within the window, thereby reducing the optimization overhead, and allowing for more focused gradient updates to accelerate convergence. Additionally, as time-steps beyond the sliding window are not involved in optimization, higher-order solvers are supported for sampling. So we present a faster variant, termed MixGRPO-Flash, which further improves training efficiency while achieving comparable performance. MixGRPO exhibits substantial gains across multiple dimensions of human preference alignment, outperforming DanceGRPO in both effectiveness and efficiency, with nearly 50% lower training time. Notably, MixGRPO-Flash further reduces training time by 71%. Codes and models are available at MixGRPO.
1 Introduction
Recent advances (Fan et al., 2023b; Xu et al., 2023; Li et al., 2025; Miao et al., 2024; Liu et al., 2025; Xue et al., 2025) in the Text-to-Image (T2I) tasks have demonstrated that flow matching models can achieve improved performance by incorporating Reinforcement Learning from Human Feedback (RLHF) (Ouyang et al., 2022) strategies during the post-training stage to maximize rewards. Specifically, methods (Liu et al., 2025; Xue et al., 2025) based on Group Relative Policy Optimization (GRPO) (Shao et al., 2024), have recently been studied, achieving optimal alignment with human preferences.
Current GRPO methods in probability flow models, e.g., Flow-GRPO (Liu et al., 2025), DanceGRPO (Xue et al., 2025), leverage Stochastic Differential Equations (SDE) sampling at every denoising step to introduce randomness in image generation, addressing reliance on stochastic exploration in RLHF. They model the denoising process as a Markov Decision Process (MDP) in a stochastic environment, using GRPO to optimize the entire state-action sequence. However, this significantly slows down the training process due to the substantial overhead introduced by the denoising iteration process. Specifically, to compute the ratio of posterior probabilities, it is essential to perform full-step sampling independently with the old policy model $\pi_{\theta_{\text{old}}}$ and the new one $\pi_{\theta}$. While DanceGRPO proposes randomly selecting a subset of denoising steps to optimize, our empirical analysis in Figure 1 demonstrates a substantial performance degradation as the subset size is reduced.
Figure 1: Performance comparison for different numbers of denoising steps optimized. The performance improvement of DanceGRPO relies on more steps optimized. MixGRPO achieves optimal performance while requiring only 4 steps.
To address these issues, we propose MixGRPO, which achieves stochastic exploration while enabling optimization over fewer denoising steps. This reduces the MDP sequence to a subset of the denoising process and uses GRPO for optimization. Specifically, we employs a mixed ODE-SDE strategy, applying SDE sampling to a subset of steps and Ordinary Differential Equations (ODE) sampling to the rest, confining randomness to SDE-based sampling steps. In this way, fewer time-steps are needed for optimization, without compromising image quality for reward computation. Besides, we introduce a sliding window strategy that moves along the denoising steps, where SDE sampling and GRPO-guided optimization are applied only within the window, while ODE sampling is used outside the window. Unlike random selection, this scheduling strategy orders optimization from high to low denoising levels, which aligns with the intuition of applying temporal discount factors to rewards in Reinforcement Learning (RL) (Pitis, 2019; Amit et al., 2020; Hu et al., 2022). MixGRPO prioritizes optimizing the initial timesteps, which involve the most significant noise removal and entail a larger exploration space (see Figure 2). Finally, we find that higher-order ODE solvers, e.g., DPMSolver++ (Lu et al., 2022b), can significantly accelerate training-time sampling, with negligible performance degradation, as there is no need for the posterior probability distribution after the sliding window.
We trained and evaluated MixGRPO by using HPS-v2.1 (Wu et al., 2023), Pick Score (Kirstain et al., 2023), ImageReward (Xu et al., 2023), and Unified Reward (Wang et al., 2025) as reward models (RMs) and metrics. We also quantified the overhead in terms of the number of function evaluations (NFE) and time consumption during training During the training process, we fine-tuned based on FLUX.1-dev (Labs, 2024) and compared performance using either a single RM or multiple RMs as guidance, assessing the results for both in-domain and out-of-domain metrics. Specifically, trained and evaluated on the HPDv2 dataset (Wu et al., 2023), MixGRPO improves the ImageReward (Xu et al., 2023) from 1.088 to 1.629, surpassing DanceGRPO’s score of 1.436, while generating images with enhanced semantic quality, aesthetics and reduced distortion. Furthermore, MixGRPO reduces the training time of DanceGRPO by nearly 50%. In addition, MixGRPO-Flash utilizes DPMSolver++ (Lu et al., 2022b) to accelerate the sampling of $\pi_{\theta_{\text{old}}}$, reducing training time by 71%.
To summarize, the key contributions of our work are outlined below:
  • We propose a mixed ODE-SDE GRPO training framework for flow matching, which alleviates the overhead bottleneck by streamlining the optimization process within the MDP.
  • We introduce a sliding window strategy to sample the denoising timesteps for model optimization, aligning with the RL intuition of transitioning from harder to easier search spaces, significantly enhancing performance.
  • Our method enables the use of higher-order ODE solvers to accelerate $\pi_{\theta_{\text{old}}}$ sampling during GRPO training, achieving significant speed improvements with comparable performance.
  • Comprehensive experiments were carried out on multiple rewards and the results demonstrate that MixGRPO achieves substantial gains on various evaluation metrics, while significantly reducing training overhead.
Figure 2: Visualization of t-SNE (Van der Maaten & Hinton, 2008) for images sampled with different strategies. Employing SDE sampling in the early stages of the denoising process results in a more discrete data distribution.
2 Related Work
2.1 RL For Image Generation
Inspired by Proximal Policy Optimization (PPO) (Schulman et al., 2017), early works (Fan & Lee, 2023; Black et al., 2023; Fan et al., 2023a; Lee et al., 2023) integrated reinforcement learning (RL) into diffusion models by optimizing the score function (Song et al., 2020b) through policy gradient methods, thereby enabling the generation of images that better align with human preferences. Subsequently,  Wallace et al. (2024) introduced offline-Direct Preference Optimization (DPO) to T2I tasks for the first time. This allows diffusion models to directly learn from human feedback and validates its effectiveness on large-scale models. Due to the tendency of offline win-lose pair data to shift the model away from its original distribution, some works (Yuan et al., 2024; Liang et al., 2025) have adopted online methods, continuously adjusting sampling trajectories through step-aware preference models during training to achieve improved performance. Recently, GRPO-based works e.g.,  Tong et al. (2025), Flow-GRPO (Liu et al., 2025) and DanceGRPO (Xue et al., 2025), have elevated RL-enhanced image generation to new heights. Specifically, Flow-GRPO (Liu et al., 2025) and DanceGRPO (Xue et al., 2025) introduced GRPO to flow matching models, allowing divergent sampling by transforming the ODE into an equivalent SDE. They also identified the overhead caused by full-step sampling within a group as a bottleneck and sought to address it by reducing or randomly selecting denoising steps. However, these approaches do not fundamentally address the issue. We hope to delve into the essence of GRPO on the probability flow and provide deeper insights through mixed sampling techniques and optimization scheduling.
2.2 Sampling Methods for Probability Flow
DDPM (Ho et al., 2020) first proposed training a probabilistic model to reverse each step of noise corruption and utilized probability flow SDE for sampling, enabling the generation of realistic images. However, such a sampling method often requires thousands of steps, resulting in significant overhead. DDIM (Song et al., 2020a) introduced deterministic sampling and proposed a probability ODE sampling approach, reducing the number of sampling steps to around 100. Subsequently, inspired by the Fokker-Planck equation (Risken & Risken, 1996), Song et al. (2020b) established a unification of SDE and ODE sampling methods from the perspective of the score function. Then, more higher-order ODE solvers were proposed e.g., DPM-Solver (Lu et al., 2022a) and DPM-Solver++ (Lu et al., 2022b), which utilize multistep methods for differential discretization. These approaches significantly reduce the number of sampling steps to around 10 while preserving accuracy. Higher-performance solvers (Zheng et al., 2023; Zhao et al., 2023) continue to be proposed; however, the gains are relatively marginal and have ultimately been replaced by the distillation method (Salimans & Ho, 2022; Yin et al., 2024). During the same period, flow matching models (Lipman et al., 2022; Esser et al., 2024) simplified and stabilized training by predicting the vector field velocity, enabling deterministic sampling with ODEs under 50 steps. Recent theoretical works (Gao et al., 2024; Albergo et al., 2023) has proven that the sampling method of flow matching is equivalent to DDIM, and demonstrated that flow matching models share the same equivalent SDE and ODE formulations as diffusion models. This provides important theoretical support and insights for our work, and we may explore interleaved sampling of SDE and ODE in probability flow models as a potential approach.
3 Method
3.1 Mixed ODE-SDE Sampling in GRPO
According to Flow-GRPO (Liu et al., 2025), the SDE sampling in flow matching can be framed as a Markov Decision Process (MDP) $\left(\mathcal{S},\mathcal{A},\rho_{0},P,\mathcal{R}\right)$ in a stochastic environment. The agent produces a trajectory during the sampling process defined as $\Gamma=(\mathbf{s}_{0},\mathbf{a}_{0},\mathbf{s}_{1},\mathbf{a}_{1},\ldots,\mathbf{s}_{T},\mathbf{a}_{T})$, which results in rewards $\mathcal{R}(\mathbf{s}_{t},\mathbf{a}_{t})\overset{\triangle}{=}R(\mathbf{x}_{T},c)$. In MixGRPO, we propose a hybrid sampling method that combines SDE and ODE. MixGRPO defines an interval $S=\left[t_{1},t_{2}\right)$, which is a subinterval of the denoising time range $\left[0,T\right]$, so $0\leq t_{1}\leq t_{2}\leq T$. During the denoising process, we use SDE sampling within the interval $S$ and ODE sampling outside, while the interval $S$ moves from 0 to $T$ throughout the training process (See Figure 2). MixGRPO restricts the agent’s stochastic exploration space to the interval $S$, shortening the sequence length of the MDP to a subset $\Gamma_{\text{MixGRPO{}}}=(\mathbf{s}_{t_{1}},\mathbf{a}_{t_{1}},\mathbf{s}_{t_{1}+1},\mathbf{a}_{t_{1}+1},\ldots,\mathbf{s}_{t_{2}})$ and requires reinforcement learning (RL) optimization only on this subset:
$\max_{\theta}\mathbb{E}_{\Gamma_{\text{MixGRPO}}\sim\pi_{\theta}}\left[\sum_{t=t_{1}}^{t_{2}-1}\left(\mathcal{R}\left(\mathbf{s}_{t},\mathbf{a}_{t}\right)-\beta D_{KL}\left(\pi\left(\cdot|\mathbf{s}_{t}\right)||\pi_{\text{ref}}\left(\cdot|\mathbf{s}_{t}\right)\right)\right)\right],$ (1)
Other settings in the MDP remain unchanged. MixGRPO reduces computational overhead while also lowering the difficulty of optimization. Next, we derive the specific sampling form and optimization objective of MixGRPO.
For a deterministic probability flow ODE (Song et al., 2020b), it takes the following form:
$\frac{\mathrm{d}\mathbf{x}_{t}}{\mathrm{d}t}=f(\mathbf{x}_{t},t)-\frac{1}{2}g^{2}(t)\nabla_{\mathbf{x}_{t}}\log{q_{t}(\mathbf{x}_{t})},\quad\mathbf{x}_{0}\sim q_{0}(\mathbf{x}_{0})$ (2)
where $f(\mathbf{x}_{t},t)$ is a vector-valued function called the drift coefficient and the $g(t)$ is a scalar function known as the diffusion coefficient. $\nabla_{\mathbf{x}_{t}}\log{q_{t}(\mathbf{x}_{t})}$ is the score function at time $t$.
According to the Fokker-Planck equation (Risken & Risken, 1996),  Song et al. (2020b) has demonstrated that Eq. (2) has the following equivalent probability flow SDE, which maintains the same marginal distribution at each time $t$:
$\frac{\mathrm{d}\mathbf{x}_{t}}{\mathrm{d}t}=f(\mathbf{x}_{t},t)-g^{2}(t)\nabla_{\mathbf{x}_{t}}\log{q_{t}(\mathbf{x}_{t})}+g(t)\frac{\mathrm{d}\mathbf{w}}{\mathrm{d}t},\quad\mathbf{x}_{0}\sim q_{0}(\mathbf{x}_{0})$ (3)
In MixGRPO, we mix ODE and SDE for sampling, with the specific form as follows:
[…]
In particular, for Flow Matching (FM) (Lipman et al., 2022), especially the Rectified Flow (RF) (Liu et al., 2022), the sampling process can be viewed as a deterministic ODE:
$\frac{\mathrm{d}\mathbf{x}_{t}}{\mathrm{d}t}=\mathbf{v}_{t},$ (5)
Eq. (5) is actually a special case of the Eq. (2) with $\mathbf{v}_{t}=f(\mathbf{x}_{t},t)-\frac{1}{2}g^{2}(t)\nabla_{\mathbf{x}_{t}}\log{q_{t}(\mathbf{x}_{t})}$. So we can derive the ODE-SDE hybrid sampling form for RF as follows:
[…]
In the RF framework, the model is used to predict the velocity field of the deterministic ODE, represented as $\mathbf{v}_{\theta}(\mathbf{x}_{t},t)=\frac{\mathrm{d}\mathbf{x}_{t}}{\mathrm{d}t}$. Following Liu et al. (2025), the score function is represented as $\nabla_{\mathbf{x}_{t}}\log{q_{t}(\mathbf{x}_{t})}=-\frac{\mathbf{x}_{t}}{t}-\frac{1-t}{t}\mathbf{v}_{\theta}(\mathbf{x}_{t},t)$. The $g(t)$ is represented as the standard deviation of the noise $g(t)=\sigma_{t}$. According to the definition of the standard Wiener process, we use $\mathrm{d}\mathbf{w}=\sqrt{\mathrm{d}t}\epsilon$, where $\epsilon\sim\mathcal{N}(0,\mathbf{I})$. Applying Euler-Maruyama discretization for SDE and Euler discretization for ODE, with $S=\{t_{1},t_{1}+1,\ldots,t_{2}-1\}$ being a set that includes all timesteps between $t_{1}$ and $t_{2}$, we build the final denoising process in MixGRPO:
[…]
The training process in MixGRPO is similar to Flow-GRPO (Liu et al., 2025) and DanceGRPO (Xue et al., 2025), but we only need to optimize for the time steps sampled within the interval $S$. The final training objective is represented as follows:
$\mathcal{J}_{\text{mixGRPO}}(\theta)$ $=\mathbb{E}_{c\sim\mathcal{C},~\{\mathbf{x}^{i}_{T}\}^{N}_{i=0}\sim\pi_{\theta_{\text{old}}}(\cdot|c)}$ (8)
$\left[\frac{1}{N}\sum^{N}_{i=1}\frac{1}{|S|}\sum_{t\in S}\text{min}\left(r^{i}_{t}(\theta)A^{i},\text{clip}\left(r^{i}_{t}(\theta),1-\varepsilon,1+\varepsilon\right)A^{i}\right)\right],$
where $r^{i}_{t}(\theta)$ is referred to as the policy ratio and $A^{i}$ is the advantage score. We set $\varepsilon=0.0001$
$r^{i}_{t}(\theta)=\frac{q_{\theta}(\mathbf{x}_{t+1}|\mathbf{x}_{t},c)}{q_{\theta_{\text{old}}}({\mathbf{x}_{t+1}|\mathbf{x}_{t},c})}\quad\text{and}\quad A^{i}=\frac{R\left(\mathbf{x}^{i}_{T},c\right)-\text{mean}\left(\{R\left(\mathbf{x}^{i}_{T},c\right)\}^{N}_{i=1}\right)}{\text{std}\left(\{R\left(\mathbf{x}^{i}_{T},c\right)\}^{N}_{i=1}\right)},$ (9)
where $R(\mathbf{x}^{i}_{0},c)$ is provided by the reward model.
It is important to note that, we have dropped the KL Loss (Shao et al., 2024). Although KL Loss can mitigate reward hacking to some extent (Liu et al., 2025), inspired by yifan123 (2025), we use mixed model sampling at test time, which can significantly address the reward hacking issue (See Appendix A).
MixGRPO reduces NFE of $\pi_{\theta}$ compared to all-timesteps optimization. However, the NFE of $\pi_{\theta_{\text{old}}}$ is not reduced, as complete inference is required to obtain the final image for reward calculation. In Section 3.3, we will introduce the use of higher-order ODE solvers, which also reduce the NFE of $\pi_{\theta_{\text{old}}}$ leading to further speedup. In summary, the mixed ODE-SDE sampling significantly reduces overhead while ensuring that the sampling does not deviate from the marginal distribution of either single ODE or SDE sampling at each timestep, due to the equivalence of the probability flow.
Algorithm 1 MixGRPO Training Process
1:initial policy model $\pi_{\theta}$; reward models $\{R_{k}\}_{k=1}^{K}$; prompt dataset $\mathcal{C}$; total sampling steps $T$; number of samples per prompt $N$;
2:sliding window $W(l)$, window size $w$, shift interval $\tau$, window stride $s$
3:Init left boundary of $W(l)$: $l\leftarrow 0$
4:for training iteration $m=1$ to $M$ do
5:  Sample batch prompts $\mathcal{C}_{b}\sim\mathcal{C}$
6:  Update old policy model: $\pi_{\theta_{\text{old}}}\leftarrow\pi_{\theta}$
7:  for each prompt $\mathbf{c}\in\mathcal{C}_{b}$ do
8:   Init the same noise $\mathbf{x_{0}}\sim\mathcal{N}(0,\mathbf{I})$
9:   for generate $i$-th image from $i=1$ todo
10:     for sampling timestep $t=0$ to $T-1$ do $\triangleright$ $\pi_{\theta_{\text{old}}}$ mixed sampling loop
11:      if $t\in W(l)$ then
12:        Use SDE Sampling to get $\mathbf{x}^{i}_{t+1}$
13:      else
14:        Use ODE Sampling to get $\mathbf{x}^{i}_{t+1}$
15:      end if
16:     end for
17:     Calculate multi-reward advantage: $A_{i}\leftarrow\sum_{k=1}^{K}\frac{R(\mathbf{x}^{i}_{T},\mathbf{c})^{i}_{k}-\mu_{k}}{\sigma_{k}}$
18:   end for
19:   for optimization timestep $t\in W(l)$ do $\triangleright$ optimize policy model $\pi_{\theta}$
20:     Update policy model via gradient ascent: $\theta\leftarrow\theta+\eta\nabla_{\theta}\mathcal{J}$
21:   end for
22:  end for
23:  if $m\bmod\tau~\text{is}~0$ then $\triangleright$ move sliding window
24:   $l\leftarrow\min(l+s,~T-w)$
25:  end if
26:end for
3.2 Sliding Window as Optimization Scheduler
Figure 3: Qualitative comparison. MixGRPO achieve superior performance in semantics, aesthetics and text-image alignment.
In fact, the interval $S$ can be non-fixed during the training process. In this section, we will introduce the sliding window to describe the movement of the interval $S$, which leads to a significant improvement in the quality of the generated images. Along the discrete denosing timesteps $\{0,1,\ldots,T-1\}$, MixGRPO defines a sliding window $W(l)$ and optimization is only employed at the timesteps within this window.
$W(l)=\{t_{l},t_{l+1},\dots,t_{l+w-1}\},\quad l\leq T-w$ (10)
where $l$ is the left boundary of the sliding window, and $w$ is a hyperparameter that represents the window size.
The left boundary $l$ of the sliding window moves as the training progresses. In the experiments, we found that the window size $w$, shift interval $\tau$, and window stride $s$ are all crucial hyperparameters. Through ablation studies (See Experiment 4.4.1), we identified the optimal settings. When the total sampling steps $T=25$, the best performance is achieved with $w=4$, $\tau=25$ and $s=1$. The detailed sliding window strategy and the MixGRPO algorithm can be referenced in Algorithm 1
Restricting the use of SDE sampling within the sliding window not only ensures the diversity of the generated images but also allows the model to concentrate on optimizing the flow within that window. Moving along the denoising direction represents the stochasticity of the probability flow from strong to weak, as illustrated in Figure 2. This is essentially a greedy strategy and is similar to the RL that assigns discount factors to process rewards (Pitis, 2019; Amit et al., 2020; Hu et al., 2022), which gives greater importance to rewards derived from a larger search space in the earlier process. The experimental results of different movement strategies in Table 4 demonstrate the validity of this intuition. Not only does the progressive strategy outperform random selection, but we also found that even when the sliding window is not moved (frozen), meaning only the earlier timesteps are optimized, MixGRPO can still yield good results, particularly in terms of ImageReward (Xu et al., 2023) and UnifiedReward (Wang et al., 2025). Based on this intuition, we also proposed an exponential decay strategy as follows, allowing $\tau$ to decrease as the window moves, enabling the model to avoid excessive optimization in smaller search spaces.
$\tau(l)=\tau_{0}\cdot\exp\left(-k\cdot\text{ReLU}\left(l-\lambda_{thr}\right)\right),$ (11)
where $\tau_{0}$ is the initial shift interval, $k$ is the decay factor, and $\lambda_{thr}$ is the threshold that controls when the decay starts. The exponential function $\exp(x)$ calculates $e^{x}$, while the Rectified Linear Unit $\text{ReLU}(x)$ is defined as $\max(0,x)$. Table 4 shows that the exponential decay strategy can achieve better results in terms of Pick Score (Kirstain et al., 2023) and ImageReward (Xu et al., 2023), likely because the model focuses on the earlier stages of denoising, which can lead to more significant high-level changes.
3.3 Trade-off Between Overhead and Performance
MixGRPO employs SDE sampling within the sliding window and ODE sampling outside of it, allowing the use of higher-order ODE solvers to accelerate GRPO training-time sampling. The timesteps that utilize ODE sampling are divided into those before and after the sliding window. The timesteps after the sliding window solely influence the reward calculation, whereas the timesteps before the window affect both the reward and contribute to cumulative errors in the policy ratio computing. Therefore, we focus exclusively on the acceleration of the timesteps after the window.
Gao et al. (2024) has demonstrated the equivalence between the ODE sampling of flow matching models (FM) and DDIM, and Section 3.1 has also shown that diffusion probabilistic models (DPM) and FM share the same ODE form during the denoising process. Therefore, the higher-order ODE solvers e.g., DPM-Solver Series (Lu et al., 2022a; b; Zheng et al., 2023), UniPC (Zhao et al., 2023) designed for DPM sampling acceleration are also applicable to FM. We have reformulated DPM-Solver++ (Lu et al., 2022b) to apply it in the FM framework for ODE sampling acceleration and released detailed derivations in Appendix B.
By applying higher-order solvers, we achieve acceleration in the sampling of $\pi_{\theta_{\text{old}}}$ during GRPO training, which is essentially a balance between overhead and performance. Excessive acceleration leads to fewer timesteps, which inevitably results in a decline in image generation quality, thereby accumulating errors in the computation of rewards. We have found in practice that the 2nd-order DPM-Solver++ is sufficient to provide significant acceleration while ensuring that the generated images align well with human preferences. Ultimately, we adopted both progressive and frozen sliding window strategies, proposing MixGRPO-Flash and MixGRPO-Flash*. A detailed description of the algorithm can be found in Appendix C. These approaches achieve a greater degree of acceleration compared to MixGRPO, while also outperforming DanceGRPO in terms of performance.
Table 1: Comparison results for overhead and performance. MixGRPO achieves the best performance across multiple metrics. MixGRPO-Flash significantly reduces training time while outperforming DanceGRPO. Bold: rank 1. Underline: rank 2. The Frozen strategy means that optimization is only employed at the initial denoising steps.
Method $\text{NFE}_{\pi_{\theta_{\text{old}}}}$ $\text{NFE}_{\pi_{\theta}}$ Iteration Time (s)$\downarrow$ Human Preference Alignment
[…]
FLUX / / / 0.313 0.227 1.088 3.370
DanceGRPO 25 14 291.284 0.356 0.233 1.436 3.397
25 4 149.978 0.334 0.225 1.335 3.374
25 4 150.059 0.333 0.229 1.235 3.325
MixGRPO 25 4 150.839 0.367 0.237 1.629 3.418
MixGRPO-Flash 16 (Avg) 4 112.372 0.358 0.236 1.528 3.407
8 4 83.278 0.357 0.232 1.624 3.402
4 Experiments
4.1 Experiment Setup
Dataset  We conduct experiments using the prompts provided by the HPDv2111https://huggingface.co/datasets/ymhao/HPDv2 dataset, which is the official dataset for the HPS-v2 benchmark (Wu et al., 2023). The training set contains 103,700 prompts; in fact, MixGRPO achieved good human preference alignment results after training one epoch with only 9,600 prompts. The test set consists of 400 prompts. The prompts are diverse, encompassing four styles: “Animation”, “Concept Art”, “Painting”, and “Photo”.
Model  Following DanceGRPO (Xue et al., 2025), we use FLUX.1 Dev (Labs, 2024) as the base model, which is an advanced text-to-image model based on flow matching.
Overhead Evaluation  For the evaluation of overhead, we use two metrics: the number of function evaluations (NFE) (Lu et al., 2022a) and the time consumption per iteration during training. The NFE is divided into $\text{NFE}_{\pi_{\theta_{\text{old}}}}$ and $\text{NFE}_{\pi_{\theta}}$. $\text{NFE}_{\pi_{\theta_{\text{old}}}}$ represents the number of forward propagation of the reference model for computing the policy ratio and generating images. $\text{NFE}_{\pi_{\theta}}$ is the number of forward propagation of the policy model solely for the policy ratio. Additionally, the average training time per GRPO iteration provides a more accurate reflection of the acceleration effect.
Performance Evaluation  We used four reward models, namely HPS-v2.1 (Wu et al., 2023), Pick Score (Kirstain et al., 2023), ImageReward (Xu et al., 2023) and Unified Reward Wang et al. (2025), both as reward guidance during training and as evaluation metrics. These metrics are all based on human preferences but emphasize different aspects. For example, ImageReward (Xu et al., 2023) highlights image-text alignment and fidelity, while Unified Reward (Wang et al., 2025) concentrates on semantics. DanceGRPO Xue et al. (2025) also demonstrates that using multiple reward models can achieve better results. To validate the robustness of MixGRPO, we also followed DanceGRPO and conducted additional comparisons using HPS-v2.1 as a single reward, and combining HPS-v2.1 (Wu et al., 2023) and CLIP Score (Radford et al., 2021) as multi-rewards.
Figure 4: Qualitative comparison with different training-time sampling steps. The performance of MixGRPO does not significantly decrease with the reduction in overhead. The Frozen strategy means that optimization is only employed at the initial denoising steps.
4.2 Implementation Details
For training-time sampling, we first perform a shift $\tilde{t}=\frac{t}{1-(\tilde{s}-1)t}$ on the uniformly distributed $t=[0,\ldots,T-1]$ and then define $\sigma_{t}=\eta\sqrt{\frac{\tilde{t}}{1-\tilde{t}}}$. We set $\tilde{s}=3$ and $\eta=0.7$ as scale. We set $T=25$ as the total sampling steps. For GRPO, the model generates 12 images for each prompt and clips the advantage to the range [5,5][-5,5][ - 5 , 5 ]. It is important to note that we use gradient accumulation over 3 steps, which means that during a single training iteration, there are $\frac{12}{3}=4$ gradient updates. For the exponential decay strategy of the sliding window in Eq. ( 11), we empirically set $\tau_{0}=20$, $k=0.1$, and $\lambda_{thr}=13$. Furthermore, when multiple reward models are jointly trained with, each reward is assigned equal weight.
For training, all experiments are conducted on 32 Nvidia GPUs with a batch size of 1 and a maximum of 300 iterations. We use AdamW (Loshchilov & Hutter, 2017) as the optimizer with a learning rate of 1e-5 and a weight decay coefficient of 0.0001. Mixed precision training is used with the bfloat16 (bf16) format, while the master weights are maintained at full precision (fp32).
4.3 Main Experiments
In the main experiment, the four human-preference-based rewards were aggregated according to advantages, as shown in Algorithm 1. We evaluated the overhead and performance of MixGRPO in comparison with DanceGRPO, with the results presented in Table 1. The official DanceGRPO uses $\text{NFE}_{\pi_{\theta}}=14$; however, for fairness, we also tested DanceGRPO with $\text{NFE}_{\pi_{\theta}}=4$. For MixGRPO-Flash, we evaluated both the progressive and frozen strategies, and to ensure fairness, we also applied the frozen strategy to DanceGRPO.
We selected multiple scene prompts to visualize the results for FLUX.1 Dev, the officially configured DanceGRPO, and MixGRPO, as shown in Figure 3. It can be observed that MixGRPO achieved the best results in terms of semantics, aesthetics, and text-image alignment. Figure 4 shows the comparison results of DanceGRPO, MixGRPO, and MixGRPO-Flash with $\text{NFE}_{\pi_{\theta}}=4$. It can be observed that under the same overhead conditions, MixGRPO achieved better results compared to DanceGRPO. Additionally, MixGRPO-Flash enables accelerated sampling of $\pi_{\theta_{\text{old}}}$, and as the overhead decreases, the quality of the generated images still maintains a strong alignment with human preferences.
Following DanceGRPO (Xue et al., 2025), we also trained and evaluated the model using a single reward model and two reward models, but on the HPDv2 dataset (Wu et al., 2023). The results (See Table 2) demonstrate that MixGRPO achieves the best performance on both in-domain and out-of-domain rewards, whether using a single or multiple reward models. The visualized results are displayed in Appendix D.
Table 2: Comparison results for in-domain and out-of-domain reward metrics. The results demonstrate that MixGRPO achieves the best performance on both in-domain and out-of-domain rewards, whether using a single or multiple reward models.
Reward Model Method In Domain Out-of-Domain
HPS-v2.1 CLIP Score Pick Score ImageReward Unified Reward
/ FLUX 0.313 0.388 0.227 1.088 3.370
HPS-v2.1 DanceGRPO 0.367 0.349 0.227 1.141 3.270
MixGRPO 0.373 0.372 0.228 1.396 3.370
HPS-v2.1 & CLIP Score DanceGRPO 0.346 0.400 0.228 1.314 3.377
MixGRPO 0.349 0.415 0.229 1.416 3.430
4.4 Ablation Experiments
4.4.1 Sliding Window Hyperparamters
As introduced in Section 3.2, the moving strategy, shift interval $\tau$, window size $w$ and window stride $s$ are all important parameters of the sliding window. We conducted ablation experiments on each of them. For the moving strategy, we compared three approaches: frozen, where the window remains stationary; random, where a random window position is selected at each iteration; and progressive, where the sliding window moves incrementally with the denoising steps. For the progressive strategy, we tested different scheduling strategies where the interval $\tau$ initially starts at 25 but changes with training iterations. As shown in Table 4, the results indicate that under the progressive strategy, either exponential decay or constant scheduling strategies are optimal. For the shift interval $\tau$, 25 is the optimal setting (See Table 4).
The number of inferences for $\pi_{\theta}$ increases with the growth of the window size $w$, leading to greater time overhead. We compared different settings of $w$, and the results are shown in Table 6. Ultimately, we selected $w=4$ as a balanced setting between overhead and performance. For the window stride $s$, we found through experimentation that $s=1$ is the optimal choice, as shown in Table 6.
Table 3: Comparison for moving strategies.
Strategy
Interval
Schedule
HPS-v2.1 Pick Score ImageReward Unified Reward
Frozen / 0.354 0.234 1.580 3.403
Random Constant 0.365 0.237 1.513 3.388
Progressive Decay (Linear) 0.365 0.235 1.566 3.382
Decay (Exp) 0.360 0.239 1.632 3.416
Constant 0.367 0.237 1.629 3.418
Table 4: Comparison for shift interval $\tau$.
$\tau$ HPS-v2.1 Pick Score ImageReward Unified Reward
15 0.366 0.237 1.509 3.403
20 0.366 0.238 1.610 3.411
25 0.367 0.237 1.629 3.418
30 0.350 0.229 1.589 3.385
Table 5: Comparison for window size $w$
$w$ $\text{NFE}_{\pi_{\theta}}$ HPS-v2.1 Pick Score ImageReward Unified Reward
2 2 0.362 0.235 1.588 3.419
4 4 0.367 0.237 1.629 3.418
6 6 0.370 0.238 1.547 3.398
Table 6: Comparison for window stride $s$
$s$ HPS-v2.1 Pick Score ImageReward Unified Reward
1 0.367 0.237 1.629 3.418
2 0.357 0.236 1.575 3.391
3 0.370 0.236 1.578 3.404
4 0.368 0.238 1.575 3.407
4.4.2 High Order ODE Solver
MixGRPO enables the possibility of accelerating ODE sampling with high-order ODE solvers by combining SDE and ODE sampling methods. We first conducted ablation experiments on the order of the solver, using DPM-Solver++ (Lu et al., 2022b) as the high-order solver with the progressive strategy. The results, as shown in Table 8, indicate that the second-order mid-point method is the optimal setting. Then, as described in Section 3.3, we compared two acceleration approaches. One is MixGRPO-Flash, which utilizes the progressive window moving strategy. The other is MixGRPO-Flash*, which employs the frozen moving strategy. They all achieve a balance between overhead and performance by reducing the number of ODE sampling steps after the sliding window. However, in practice, MixGRPO-Flash requires the window to move throughout the training process, which results in a shorter ODE portion being accelerated. Consequently, the acceleration effect of MixGRPO-Flash, on average, is not as pronounced as that of MixGRPO-Flash*.
Table 7: Comparison of performance across different-order solvers. The second-order Midpoint method achieves the best performance.
Order Solver Type HPS-v2.1 Pick Score ImageReward Unified Reward
1 / 0.367 0.236 1.570 3.403
2 Midpoint 0.358 0.237 1.578 3.407
Heun 0.362 0.233 1.488 3.399
3 / 0.359 0.234 1.512 3.387
Table 8: Comparison of different sampling steps for $\pi_{\theta_{\text{old}}}$. MixGRPO-Flash* achieves good performance even with few steps.
Method Sampling Overhead Human Preference Alignment
$\text{NFE}_{\pi_{\theta_{\text{old}}}}$ Time per Image (s) HPS-v2.1 Pick Score ImageReward Unified Reward
DanceGRPO 25 9.301 0.334 0.225 1.335 3.374
MixGRPO-Flash 19 (Avg) 7.343 0.357 0.236 1.564 3.394
16 (Avg) 6.426 0.362 0.237 1.578 3.407
13 (Avg) 5.453 0.344 0.229 1.447 3.363
MixGRPO-Flash* 12 4.859 0.353 0.230 1.588 3.396
10 4.214 0.359 0.234 1.548 3.430
8 3.789 0.357 0.232 1.624 3.402
5 Conclusion
Although GRPO (Shao et al., 2024) has seen significant success in the language modality, it is still in the early stages of progress in vision (Tong et al., 2025; Xue et al., 2025; Liu et al., 2025). Existing flow-matching based GRPO faces challenges such as low sampling efficiency and slow training. To address these issues, we proposed MixGRPO, a novel training framework that combines SDE and ODE sampling. This hybrid approach allows for focused optimization on the SDE sampling flow component, reducing complexity while ensuring accurate reward computation. Inspired by the decay factor in reinforcement learning (Hu et al., 2022), we introduce a sliding window strategy for scheduling the optimized denoising steps. Experimental results confirm the effectiveness of our approach in both single-reward and multi-reward settings. Additionally, MixGRPO decouples the denoising stages for optimization and reward computation, enabling acceleration of the latter with high-order solvers. We further propose MixGRPO-Flash, which balances overhead and performance. We hope MixGRPO will inspire deeper research into post-training for image generation, contributing to the advancement of Artificial General Intelligence (AGI).
Appendix A Hybrid Inference for Solving Reward Hacking
To address the reward hacking issue, we employ hybrid inference to sample with both the raw and post-trained models, and introduce the hybrid percent $p_{\text{mix}}$. This means that the initial $p_{\text{mix}}T$ denoising steps are sampled by the model trained with GRPO, while the remaining denoising process is finished by the original FLUX model (Black et al., 2023). Table 9 and Figure 5 respectively illustrate the changes in performance and images as $p_{\text{mix}}$ increases under the multi-rewards training scenario. The experimental results demonstrate that $p_{\text{mix}}=80\%$ is an optimal empirical value that effectively mitigates hacking while maximizing alignment with human preferences.
Table 9: Comparison with different hybrid inference percentages
$p_{\text{mix}}$ HPS-v2.1 Pick Score ImageReward Unified Reward
0% 0.313 0.226 1.089 3.369
20% 0.342 0.233 1.372 3.386
40% 0.356 0.235 1.539 3.395
60% 0.362 0.236 1.598 3.407
80% 0.366 0.238 1.610 3.411
100% 0.369 0.238 1.607 3.378
Figure 5: Qualitative comparison with different hybrid inference percentages
Appendix B DPM-Solver++ for Recitified Flow
The DPM-Solver++ algorithm (Lu et al., 2022b) is originally designed for the $\mathbf{x}_{0}\text{-}prediction$ diffusion model (Rombach et al., 2022), where the model outputs the denoised feature $\mathbf{x}_{0}$ based on the noisy feature $\mathbf{x}_{t}$, the time condition $t$, and the text condition $c$. According to the definition of Rectified Flow (RF) (Liu et al., 2022), there is the following transfer equation:
$\mathbf{x}_{t}=t\mathbf{x}_{1}+(1-t)\mathbf{x}_{0}$ (12)
According to the theory of stochastic interpolation (Albergo et al., 2023), RF effectively approximates $\mathbf{x}_{1}-\mathbf{x}_{0}$ by modeling $\mathbf{v}_{t}$:
$\mathbf{v}_{t}=\mathbf{x}_{1}-\mathbf{x}_{0}$ (13)
Based on Eq. (12) and Eq. (13), we obtain the following relationship:
$\mathbf{x}_{0}=\mathbf{x}_{t}-\mathbf{v}_{t}t$ (14)
By using a neural network for approximation, we establish the relationship between RF and the $\mathbf{x}_{0}\text{-}prediction$ model:
$\mathbf{x}_{\theta}=\mathbf{x}_{t}-\mathbf{v}_{\theta}t$ (15)
Taking the multistep second-order solver as an example (see Algorithm 2 in  (Lu et al., 2022b)), we derive the corrected $\mathbf{x}_{\theta}$ for the RF sampling process as $\mathbf{D}_{t}$:
$\mathbf{D}_{t}\leftarrow\left(1+\frac{h_{t}}{2h_{t-1}}\right)\left(\mathbf{x}_{t-1}-\mathbf{v}_{\theta}(\mathbf{x}_{t-1},t-1,c)\frac{t-1}{T}\right)$ (16)
$-\frac{h_{t}}{2h_{t-1}}\left(\mathbf{x}_{t-2}-\mathbf{v}_{\theta}(\mathbf{x}_{t-2},t-2,c)\frac{t-2}{T}\right)$
where $T$ is the total sampling steps and $h_{t}=\lambda_{t}-\lambda_{t-1}$. $\lambda_{t}$ is the log-signal-to-noise-ratio (log-SNR) and is defined in RF as $\lambda_{t}:=\log\left(\frac{1-t}{t}\right)$.
Based on the exact discretization formula for the probability flow ODE proposed in DPM-Solver++ (Eq. (9) in (Lu et al., 2022b)), we can derive the final transfer equation:
$\mathbf{x}_{t}\leftarrow\frac{t}{t-1}\mathbf{x}_{t-1}-(1-t)\left(e^{-h_{t}}-1\right)\mathbf{D}_{t},\quad 1\leq t<T$ (17)
Appendix C MixGRPO-Flash Algorithm
MixGRPO-Flash accelerates the ODE sampling that does not contribute to the calculation of the policy ratio after the sliding window by using DPM-Solver++ in the Eq. (17). We introduce a compression rate $\tilde{r}$ such that the ODE sampling after the window only requires $(T-l-w)\tilde{r}$ time steps. And the total time-steps is $\tilde{T}=l+w+(T-l-w)\tilde{r}$ The final algorithm is as follows:
Algorithm 2 MixGRPO-Flash Training Process
1:initial policy model $\pi_{\theta}$; reward models $\{R_{k}\}_{k=1}^{K}$; prompt dataset $\mathcal{C}$; total sampling steps $\tilde{T}$; number of samples per prompt $N$; ODE compression rate $\tilde{r}$
2:sliding window $W(l)$, window size $w$, shift interval $\tau$, window stride $s$
3:Init left boundary of $W(l)$: $l\leftarrow 0$
4:for training iteration $m=1$ to $M$ do
5:  Sample batch prompts $\mathcal{C}_{b}\sim\mathcal{C}$
6:  Update old policy model: $\pi_{\theta_{\text{old}}}\leftarrow\pi_{\theta}$
7:  for each prompt $\mathbf{c}\in\mathcal{C}_{b}$ do
8:   Init the same noise $\mathbf{x_{0}}\sim\mathcal{N}(0,\mathbf{I})$
9:   for generate $i$-th image from $i=1$ todo
10:     for sampling timestep $t=0$ to $\tilde{T}-1$ do
11:      if $t<l$ then
12:        Use first-order ODE sampling to get $\mathbf{x}^{i}_{t+1}$
13:      else if  $l\leq t<l+w$ then
14:        Use SDE sampling to get $\mathbf{x}^{i}_{t+1}$
15:      $\triangleright$ DPM-Solver++
16:        Use higher-order ODE sampling to get $\mathbf{x}^{i}_{t+1}$
17:      end if
18:     end for
19:     Calculate multi-reward advantage: $A_{i}\leftarrow\sum_{k=1}^{K}\frac{R(\mathbf{x}^{i}_{\tilde{T}},\mathbf{c})^{i}_{k}-\mu_{k}}{\sigma_{k}}$
20:   end for
21:   for optimization timestep $t\in W(l)$ do $\triangleright$ optimize policy model $\pi_{\theta}$
22:     Update policy model via gradient ascent: $\theta\leftarrow\theta+\eta\nabla_{\theta}\mathcal{J}$
23:   end for
24:  end for
25:  if use MixGRPO-Flash* then $\triangleright$ move sliding window
26:   $l\leftarrow 0$
27:  else
28:   if $m\bmod\tau~\text{is}~0$ then
29:     $l\leftarrow\min(l+s,~T-w)$
30:   end if
31:  end if
32:end for
Note that when using MixGRPO-Flash*, the frozen strategy is applied, with the left boundary of the sliding window $l\equiv 0$. The theoretical speedup of the training-time sampling can be described as follows:
$S=\frac{T}{w+(T-w)\tilde{r}}$ (18)
For MixGRPO-Flash, since the sliding window moves according to the progressive strategy during training, the average speedup can be expressed in the following form:
$S=\frac{T}{\mathbb{E}_{l}\left(w+l+\lceil(T-w-l)\tilde{r}\rceil\right)}<\frac{T}{w+(T-w)\tilde{r}}$ (19)
Appendix D More visualized results
Figure 6: Comparison of the visualization results of FLUX, DanceGRPO, and MixGRPO under HPS-v2.1 as the reward model.
Figure 7: Comparison of the visualization results of FLUX, DanceGRPO, and MixGRPO under HPS-v2.1 and CLIP Score as the reward models.

### v2

@@
 Abstract
-Although GRPO substantially enhances flow matching models in human preference alignment of image generation, methods such as FlowGRPO still exhibit inefficiency due to the necessity of sampling and optimizing over all denoising steps specified by the Markov Decision Process (MDP).
+Although GRPO substantially enhances flow matching models in human preference alignment of image generation, methods such as DanceGRPO still exhibit inefficiency due to the necessity of sampling and optimizing over all denoising steps specified by the Markov Decision Process (MDP).
 In this paper, we propose MixGRPO, a novel framework that leverages the flexibility of mixed sampling strategies through the integration of stochastic differential equations (SDE) and ordinary differential equations (ODE).
@@
 1 Introduction
-Recent advances (Fan et al., 2023b; Xu et al., 2023; Li et al., 2025; Miao et al., 2024; Liu et al., 2025; Xue et al., 2025) in the Text-to-Image (T2I) tasks have demonstrated that flow matching models can achieve improved performance by incorporating Reinforcement Learning from Human Feedback (RLHF) (Ouyang et al., 2022) strategies during the post-training stage to maximize rewards.
-Specifically, methods (Liu et al., 2025; Xue et al., 2025) based on Group Relative Policy Optimization (GRPO) (Shao et al., 2024), have recently been studied, achieving optimal alignment with human preferences.
-Current GRPO methods in probability flow models, e.g., Flow-GRPO (Liu et al., 2025), DanceGRPO (Xue et al., 2025), leverage Stochastic Differential Equations (SDE) sampling at every denoising step to introduce randomness in image generation, addressing reliance on stochastic exploration in RLHF.
-They model the denoising process as a Markov Decision Process (MDP) in a stochastic environment, using GRPO to optimize the entire state-action sequence.
-However, this significantly slows down the training process due to the substantial overhead introduced by the denoising iteration process.
-Specifically, to compute the ratio of posterior probabilities, it is essential to perform full-step sampling independently with the old policy model $\pi_{\theta_{\text{old}}}$ and the new one $\pi_{\theta}$.
-While DanceGRPO proposes randomly selecting a subset of denoising steps to optimize, our empirical analysis in Figure 1 demonstrates a substantial performance degradation as the subset size is reduced.
+Recent advances (Liu et al., 2025a; Xue et al., 2025; Liu et al., 2025b; Li et al., 2025; Xu et al., 2023) in Text-to-Image (T2I) tasks have demonstrated that probability flow models can achieve improved performance by incorporating Reinforcement Learning from Human Feedback (RLHF) (Ouyang et al., 2022) strategies during the post-training stage to maximize rewards.
+Specifically, methods (Liu et al., 2025a; Xue et al., 2025) based on Group Relative Policy Optimization (GRPO) (Shao et al., 2024), have recently been studied, achieving optimal alignment with human preferences.
+Current GRPO methods in probability flow models, e.g., Flow-GRPO (Liu et al., 2025a), DanceGRPO (Xue et al., 2025), leverage Stochastic Differential Equations (SDE) sampling at every denoising step to introduce randomness, addressing reliance on stochastic exploration in RLHF.
+They model the entire denoising process as a Markov Decision Process (MDP) in a stochastic environment, using GRPO to optimize the state-action sequence.
+However, the need to optimize all denoising steps not only increases overhead but also leads to inconsistent gradient descent, resulting in inefficient training.
+Specifically, to compute the policy ratio, it is essential to perform full-step sampling independently with the old policy model $\pi_{\theta_{\text{old}}}$ and new one $\pi_{\theta}$.
+While DanceGRPO proposes to randomly select a subset of steps at a fixed ratio to optimize, our empirical analysis in Figure 1 demonstrates a substantial performance degradation as the subset size is reduced.
 Figure 1: Performance comparison for different numbers of denoising steps optimized.
@@
 MixGRPO achieves optimal performance while requiring only 4 steps.
-To address these issues, we propose MixGRPO, which achieves stochastic exploration while enabling optimization over fewer denoising steps.
-This reduces the MDP sequence to a subset of the denoising process and uses GRPO for optimization.
-Specifically, we employs a mixed ODE-SDE strategy, applying SDE sampling to a subset of steps and Ordinary Differential Equations (ODE) sampling to the rest, confining randomness to SDE-based sampling steps.
-In this way, fewer time-steps are needed for optimization, without compromising image quality for reward computation.
-Besides, we introduce a sliding window strategy that moves along the denoising steps, where SDE sampling and GRPO-guided optimization are applied only within the window, while ODE sampling is used outside the window.
-Unlike random selection, this scheduling strategy orders optimization from high to low denoising levels, which aligns with the intuition of applying temporal discount factors to rewards in Reinforcement Learning (RL) (Pitis, 2019; Amit et al., 2020; Hu et al., 2022).
+To address these issues, we propose MixGRPO, which achieves more efficient stochastic exploration while enabling optimization over fewer denoising steps.
+Specifically, we employ a mixed ODE-SDE strategy, applying SDE sampling to a denoising sub-interval and Ordinary Differential Equations (ODE) sampling to the rest, confining randomness to the SDE interval.
+In this way, fewer time-steps are needed for GRPO optimization, without compromising image quality for reward computation.
+Besides, we introduce a sliding window strategy for the SDE interval that moves along the denoising steps as the training process progresses.
+Unlike random selection, this scheduling strategy orders optimization from high to low denoising levels, which aligns with the intuition of applying temporal discount factors to rewards in Reinforcement Learning (RL) (Pitis, 2019; Amit et al., 2020; Hu et al., 2022b).
 MixGRPO prioritizes optimizing the initial timesteps, which involve the most significant noise removal and entail a larger exploration space (see Figure 2).
+As training progresses, MixGRPO continuously moves the SDE window to narrow the exploration space and achieve global optimization.
 Finally, we find that higher-order ODE solvers, e.g., DPMSolver++ (Lu et al., 2022b), can significantly accelerate training-time sampling, with negligible performance degradation, as there is no need for the posterior probability distribution after the sliding window.
 We trained and evaluated MixGRPO by using HPS-v2.1 (Wu et al., 2023), Pick Score (Kirstain et al., 2023), ImageReward (Xu et al., 2023), and Unified Reward (Wang et al., 2025) as reward models (RMs) and metrics.
-We also quantified the overhead in terms of the number of function evaluations (NFE) and time consumption during training During the training process, we fine-tuned based on FLUX.1-dev (Labs, 2024) and compared performance using either a single RM or multiple RMs as guidance, assessing the results for both in-domain and out-of-domain metrics.
-Specifically, trained and evaluated on the HPDv2 dataset (Wu et al., 2023), MixGRPO improves the ImageReward (Xu et al., 2023) from 1.088 to 1.629, surpassing DanceGRPO’s score of 1.436, while generating images with enhanced semantic quality, aesthetics and reduced distortion.
+We also quantified the overhead in terms of the number of function evaluations (NFE) and time consumption during training.
+During the training process, we fine-tuned based on FLUX.1-dev (Labs, 2024) and compared performance using either a single RM or multiple RMs as guidance, assessing the results for both in-domain and out-of-domain metrics.
+Specifically, trained and evaluated on the HPDv2 dataset (Wu et al., 2023), MixGRPO outperforms DanceGRPO across all metrics, particularly improving the ImageReward (Xu et al., 2023) from 1.088 to 1.629, surpassing DanceGRPO’s score of 1.436, while generating images with enhanced semantic quality, aesthetics, and reduced distortion.
 Furthermore, MixGRPO reduces the training time of DanceGRPO by nearly 50%.
@@
 To summarize, the key contributions of our work are outlined below:
-  • We propose a mixed ODE-SDE GRPO training framework for flow matching, which alleviates the overhead bottleneck by streamlining the optimization process within the MDP.
+  • We propose a mixed ODE-SDE GRPO training framework for flow-based models, which alleviates the overhead bottleneck by streamlining the optimization process within the MDP.
   • We introduce a sliding window strategy to sample the denoising timesteps for model optimization, aligning with the RL intuition of transitioning from harder to easier search spaces, significantly enhancing performance.
-  • Our method enables the use of higher-order ODE solvers to accelerate $\pi_{\theta_{\text{old}}}$ sampling during GRPO training, achieving significant speed improvements with comparable performance.
+  • Our method enables the use of higher-order ODE solvers to accelerate $\pi_{\theta_{\text{old}}}$ sampling during GRPO training, achieving more significant speed improvements with comparable performance.
   • Comprehensive experiments were carried out on multiple rewards and the results demonstrate that MixGRPO achieves substantial gains on various evaluation metrics, while significantly reducing training overhead.
@@
 2.1 RL For Image Generation
-Inspired by Proximal Policy Optimization (PPO) (Schulman et al., 2017), early works (Fan & Lee, 2023; Black et al., 2023; Fan et al., 2023a; Lee et al., 2023) integrated reinforcement learning (RL) into diffusion models by optimizing the score function (Song et al., 2020b) through policy gradient methods, thereby enabling the generation of images that better align with human preferences.
+Inspired by Proximal Policy Optimization (PPO) (Schulman et al., 2017), early works (Fan & Lee, 2023; Black et al., 2023; Fan et al., 2023; Lee et al., 2023) integrated reinforcement learning (RL) into diffusion models by optimizing the score function (Song et al., 2020b) through policy gradient methods, thereby enabling the generation of images that better align with human preferences.
 Subsequently,  Wallace et al.
@@
 Recently, GRPO-based works e.g.,  Tong et al.
-(2025), Flow-GRPO (Liu et al., 2025) and DanceGRPO (Xue et al., 2025), have elevated RL-enhanced image generation to new heights.
-Specifically, Flow-GRPO (Liu et al., 2025) and DanceGRPO (Xue et al., 2025) introduced GRPO to flow matching models, allowing divergent sampling by transforming the ODE into an equivalent SDE.
+(2025), Flow-GRPO (Liu et al., 2025a) and DanceGRPO (Xue et al., 2025), have elevated RL-enhanced image generation to new heights.
+Specifically, Flow-GRPO (Liu et al., 2025a) and DanceGRPO (Xue et al., 2025) introduced GRPO to flow matching models, allowing divergent sampling by transforming the ODE into an equivalent SDE.
 They also identified the overhead caused by full-step sampling within a group as a bottleneck and sought to address it by reducing or randomly selecting denoising steps.
@@
 DDPM (Ho et al., 2020) first proposed training a probabilistic model to reverse each step of noise corruption and utilized probability flow SDE for sampling, enabling the generation of realistic images.
-However, such a sampling method often requires thousands of steps, resulting in significant overhead.
+However, this often requires thousands of steps, resulting in significant overhead.
 DDIM (Song et al., 2020a) introduced deterministic sampling and proposed a probability ODE sampling approach, reducing the number of sampling steps to around 100.
@@
 3.1 Mixed ODE-SDE Sampling in GRPO
-According to Flow-GRPO (Liu et al., 2025), the SDE sampling in flow matching can be framed as a Markov Decision Process (MDP) $\left(\mathcal{S},\mathcal{A},\rho_{0},P,\mathcal{R}\right)$ in a stochastic environment.
-The agent produces a trajectory during the sampling process defined as $\Gamma=(\mathbf{s}_{0},\mathbf{a}_{0},\mathbf{s}_{1},\mathbf{a}_{1},\ldots,\mathbf{s}_{T},\mathbf{a}_{T})$, which results in rewards $\mathcal{R}(\mathbf{s}_{t},\mathbf{a}_{t})\overset{\triangle}{=}R(\mathbf{x}_{T},c)$.
+According to Flow-GRPO (Liu et al., 2025a), the SDE sampling in flow matching can be framed as a Markov Decision Process (MDP) $\left(\mathcal{S},\mathcal{A},\rho_{0},P,\mathcal{R}\right)$ in a stochastic environment.
+The agent produces a trajectory during the discrete sampling process defined as $\Gamma=(\mathbf{s}_{0},\mathbf{a}_{0},\mathbf{s}_{1},\mathbf{a}_{1},\ldots,\mathbf{s}_{T},\mathbf{a}_{T})$, where the reward is provided only at the final step by the reward model, specifically $\mathcal{R}(\mathbf{s}_{i},\mathbf{a}_{i})\overset{\triangle}{=}R(\mathbf{x}_{T},c)$ if $i=T$, and 0 otherwise.
 In MixGRPO, we propose a hybrid sampling method that combines SDE and ODE.
-MixGRPO defines an interval $S=\left[t_{1},t_{2}\right)$, which is a subinterval of the denoising time range $\left[0,T\right]$, so $0\leq t_{1}\leq t_{2}\leq T$.
-During the denoising process, we use SDE sampling within the interval $S$ and ODE sampling outside, while the interval $S$ moves from 0 to $T$ throughout the training process (See Figure 2).
-MixGRPO restricts the agent’s stochastic exploration space to the interval $S$, shortening the sequence length of the MDP to a subset $\Gamma_{\text{MixGRPO{}}}=(\mathbf{s}_{t_{1}},\mathbf{a}_{t_{1}},\mathbf{s}_{t_{1}+1},\mathbf{a}_{t_{1}+1},\ldots,\mathbf{s}_{t_{2}})$ and requires reinforcement learning (RL) optimization only on this subset:
-$\max_{\theta}\mathbb{E}_{\Gamma_{\text{MixGRPO}}\sim\pi_{\theta}}\left[\sum_{t=t_{1}}^{t_{2}-1}\left(\mathcal{R}\left(\mathbf{s}_{t},\mathbf{a}_{t}\right)-\beta D_{KL}\left(\pi\left(\cdot|\mathbf{s}_{t}\right)||\pi_{\text{ref}}\left(\cdot|\mathbf{s}_{t}\right)\right)\right)\right],$ (1)
-Other settings in the MDP remain unchanged.
+MixGRPO defines a time interval $S=[t_{l},t_{r})\in[0,1)$, which corresponds to a subinterval of denoising timesteps, such that $0\leq l<r\leq T$ and $t_{i}=\frac{i}{T}$.
+We use SDE sampling within the interval $S$ and ODE sampling outside, while $S$ shifts along the denoising direction throughout the training process (See Figure 2).
+MixGRPO restricts the agent’s stochastic exploration space to the interval $S$, shortening the sequence length of the MDP to a subset $\Gamma_{\text{MixGRPO{}}}=(\mathbf{s}_{l},\mathbf{a}_{l},\mathbf{s}_{l+1},\mathbf{a}_{l+1},\ldots,\mathbf{s}_{r},\mathbf{a}_{r})$ and requires reinforcement learning (RL) optimization only on this subset:
+$\max_{\theta}\mathbb{E}_{\Gamma_{\text{MixGRPO}}\sim\pi_{\theta}}\left[\sum_{t\in S}\big(\mathcal{R}\left(\mathbf{s}_{t},\mathbf{a}_{t}\right)-\beta D_{KL}\left(\pi\left(\cdot|\mathbf{s}_{t}\right)||\pi_{\text{ref}}\left(\cdot|\mathbf{s}_{t}\right)\right)\big)\right],$ (1)
+where $\mathcal{R}(\mathbf{s}_{i},\mathbf{a}_{i})\overset{\triangle}{=}R(\mathbf{x}_{T},c)$ if $i=r$, and 0 otherwise.
 MixGRPO reduces computational overhead while also lowering the difficulty of optimization.
 Next, we derive the specific sampling form and optimization objective of MixGRPO.
-For a deterministic probability flow ODE (Song et al., 2020b), it takes the following form:
-$\frac{\mathrm{d}\mathbf{x}_{t}}{\mathrm{d}t}=f(\mathbf{x}_{t},t)-\frac{1}{2}g^{2}(t)\nabla_{\mathbf{x}_{t}}\log{q_{t}(\mathbf{x}_{t})},\quad\mathbf{x}_{0}\sim q_{0}(\mathbf{x}_{0})$ (2)
-where $f(\mathbf{x}_{t},t)$ is a vector-valued function called the drift coefficient and the $g(t)$ is a scalar function known as the diffusion coefficient.
+For a deterministic reverse probability flow ODE (Song et al., 2020b), it takes the following form:
+$\frac{\mathrm{d}\mathbf{x}_{t}}{\mathrm{d}t}=f(\mathbf{x}_{t},t)-\frac{1}{2}g^{2}(t)\nabla_{\mathbf{x}_{t}}\log{q_{t}(\mathbf{x}_{t})},\quad\mathbf{x}_{0}\sim q_{0}(\mathbf{x}_{0}),$ (2)
+where $q_{t}(\mathbf{x}_{t})$ represents the evolution process of the reverse probability distribution from 0 to $T$.
 $\nabla_{\mathbf{x}_{t}}\log{q_{t}(\mathbf{x}_{t})}$ is the score function at time $t$.
-According to the Fokker-Planck equation (Risken & Risken, 1996),  Song et al.
+According to the Fokker-Planck equation (Risken & Risken, 1996; Øksendal, 2003),  Song et al.
 (2020b) has demonstrated that Eq.
 (2) has the following equivalent probability flow SDE, which maintains the same marginal distribution at each time $t$:
-$\frac{\mathrm{d}\mathbf{x}_{t}}{\mathrm{d}t}=f(\mathbf{x}_{t},t)-g^{2}(t)\nabla_{\mathbf{x}_{t}}\log{q_{t}(\mathbf{x}_{t})}+g(t)\frac{\mathrm{d}\mathbf{w}}{\mathrm{d}t},\quad\mathbf{x}_{0}\sim q_{0}(\mathbf{x}_{0})$ (3)
-In MixGRPO, we mix ODE and SDE for sampling, with the specific form as follows:
+$\frac{\mathrm{d}\mathbf{x}_{t}}{\mathrm{d}t}=f(\mathbf{x}_{t},t)-g^{2}(t)\nabla_{\mathbf{x}_{t}}\log{q_{t}(\mathbf{x}_{t})}+g(t)\frac{\mathrm{d}\mathbf{w}}{\mathrm{d}t},\quad\mathbf{x}_{0}\sim q_{0}(\mathbf{x}_{0}).$ (3)
+In MixGRPO, we mix ODE and SDE for sampling, which has the same convergence as using only ODE sampling (a detailed proof in Appendix A). The specific form is as follows:
 […]
 In particular, for Flow Matching (FM) (Lipman et al., 2022), especially the Rectified Flow (RF) (Liu et al., 2022), the sampling process can be viewed as a deterministic ODE:
-$\frac{\mathrm{d}\mathbf{x}_{t}}{\mathrm{d}t}=\mathbf{v}_{t},$ (5)
+$\frac{\mathrm{d}\mathbf{x}_{t}}{\mathrm{d}t}=\mathbf{v}_{t}.$ (5)
 Eq.
@@
 Following Liu et al.
-(2025), the score function is represented as $\nabla_{\mathbf{x}_{t}}\log{q_{t}(\mathbf{x}_{t})}=-\frac{\mathbf{x}_{t}}{t}-\frac{1-t}{t}\mathbf{v}_{\theta}(\mathbf{x}_{t},t)$.
+(2025a), the score function is represented as $\nabla_{\mathbf{x}_{t}}\log{q_{t}(\mathbf{x}_{t})}=-\frac{\mathbf{x}_{t}}{t}-\frac{1-t}{t}\mathbf{v}_{\theta}(\mathbf{x}_{t},t)$.
 The $g(t)$ is represented as the standard deviation of the noise $g(t)=\sigma_{t}$.
 According to the definition of the standard Wiener process, we use $\mathrm{d}\mathbf{w}=\sqrt{\mathrm{d}t}\epsilon$, where $\epsilon\sim\mathcal{N}(0,\mathbf{I})$.
-Applying Euler-Maruyama discretization for SDE and Euler discretization for ODE, with $S=\{t_{1},t_{1}+1,\ldots,t_{2}-1\}$ being a set that includes all timesteps between $t_{1}$ and $t_{2}$, we build the final denoising process in MixGRPO:
+Applying Euler-Maruyama discretization for SDE and Euler discretization for ODE, we build the final denoising process in MixGRPO:
 […]
-The training process in MixGRPO is similar to Flow-GRPO (Liu et al., 2025) and DanceGRPO (Xue et al., 2025), but we only need to optimize for the time steps sampled within the interval $S$.
-The final training objective is represented as follows:
+According to Eq. (1), we only need to optimize GRPO (Shao et al., 2024) at the time within the interval $S$ for $N$ samples in the group. The final training objective is represented as follows:
 $\mathcal{J}_{\text{MixGRPO}}(\theta)$ $=\mathbb{E}_{c\sim\mathcal{C},~\{\mathbf{x}^{i}_{T}\}^{N}_{i=0}\sim\pi_{\theta_{\text{old}}}(\cdot|c)}$ (8)
 $\left[\frac{1}{N}\sum^{N}_{i=1}\frac{1}{|S|}\sum_{t\in S}\text{min}\left(r^{i}_{t}(\theta)A^{i},\text{clip}\left(r^{i}_{t}(\theta),1-\varepsilon,1+\varepsilon\right)A^{i}\right)\right],$
-where $r^{i}_{t}(\theta)$ is referred to as the policy ratio and $A^{i}$ is the advantage score. We set $\varepsilon=0.0001$
-$r^{i}_{t}(\theta)=\frac{q_{\theta}(\mathbf{x}_{t+1}|\mathbf{x}_{t},c)}{q_{\theta_{\text{old}}}({\mathbf{x}_{t+1}|\mathbf{x}_{t},c})}\quad\text{and}\quad A^{i}=\frac{R\left(\mathbf{x}^{i}_{T},c\right)-\text{mean}\left(\{R\left(\mathbf{x}^{i}_{T},c\right)\}^{N}_{i=1}\right)}{\text{std}\left(\{R\left(\mathbf{x}^{i}_{T},c\right)\}^{N}_{i=1}\right)},$ (9)
-where $R(\mathbf{x}^{i}_{0},c)$ is provided by the reward model.
-It is important to note that, we have dropped the KL Loss (Shao et al., 2024).
-Although KL Loss can mitigate reward hacking to some extent (Liu et al., 2025), inspired by yifan123 (2025), we use mixed model sampling at test time, which can significantly address the reward hacking issue (See Appendix A).
+where $r^{i}_{t}(\theta)$ is referred to as the policy ratio and $A^{i}$ is the advantage score. We set $\varepsilon=0.0001$,
+$r^{i}_{t}(\theta)=\frac{q_{\theta}(\mathbf{x}_{t+\Delta t}|\mathbf{x}_{t},c)}{q_{\theta_{\text{old}}}({\mathbf{x}_{t+\Delta t}|\mathbf{x}_{t},c})}\quad\text{and}\quad A^{i}=\frac{R\left(\mathbf{x}^{i}_{T},c\right)-\text{mean}\left(\{R\left(\mathbf{x}^{i}_{T},c\right)\}^{N}_{i=1}\right)}{\text{std}\left(\{R\left(\mathbf{x}^{i}_{T},c\right)\}^{N}_{i=1}\right)},$ (9)
+It is important to note that, we have dropped the KL Loss.
+Although KL Loss can mitigate reward hacking to some extent (Liu et al., 2025a), inspired by yifan123 (2025), we use hybrid inference at test time, which can significantly address the reward hacking issue (See Appendix B).
 MixGRPO reduces NFE of $\pi_{\theta}$ compared to all-timesteps optimization.
@@
 In Section 3.3, we will introduce the use of higher-order ODE solvers, which also reduce the NFE of $\pi_{\theta_{\text{old}}}$ leading to further speedup.
-In summary, the mixed ODE-SDE sampling significantly reduces overhead while ensuring that the sampling does not deviate from the marginal distribution of either single ODE or SDE sampling at each timestep, due to the equivalence of the probability flow.
-Algorithm 1 MixGRPO Training Process
-1:initial policy model $\pi_{\theta}$; reward models $\{R_{k}\}_{k=1}^{K}$; prompt dataset $\mathcal{C}$; total sampling steps $T$; number of samples per prompt $N$;
-2:sliding window $W(l)$, window size $w$, shift interval $\tau$, window stride $s$
-3:Init left boundary of $W(l)$: $l\leftarrow 0$
-4:for training iteration $m=1$ to $M$ do
-5:  Sample batch prompts $\mathcal{C}_{b}\sim\mathcal{C}$
-6:  Update old policy model: $\pi_{\theta_{\text{old}}}\leftarrow\pi_{\theta}$
-7:  for each prompt $\mathbf{c}\in\mathcal{C}_{b}$ do
-8:   Init the same noise $\mathbf{x_{0}}\sim\mathcal{N}(0,\mathbf{I})$
-9:   for generate $i$-th image from $i=1$ todo
-10:     for sampling timestep $t=0$ to $T-1$ do $\triangleright$ $\pi_{\theta_{\text{old}}}$ mixed sampling loop
-11:      if $t\in W(l)$ then
-12:        Use SDE Sampling to get $\mathbf{x}^{i}_{t+1}$
-13:      else
-14:        Use ODE Sampling to get $\mathbf{x}^{i}_{t+1}$
-15:      end if
-16:     end for
-17:     Calculate multi-reward advantage: $A_{i}\leftarrow\sum_{k=1}^{K}\frac{R(\mathbf{x}^{i}_{T},\mathbf{c})^{i}_{k}-\mu_{k}}{\sigma_{k}}$
-18:   end for
-19:   for optimization timestep $t\in W(l)$ do $\triangleright$ optimize policy model $\pi_{\theta}$
-20:     Update policy model via gradient ascent: $\theta\leftarrow\theta+\eta\nabla_{\theta}\mathcal{J}$
-21:   end for
-22:  end for
-23:  if $m\bmod\tau~\text{is}~0$ then $\triangleright$ move sliding window
-24:   $l\leftarrow\min(l+s,~T-w)$
-25:  end if
-26:end for
+In summary, the mixed ODE-SDE sampling significantly reduces overhead while ensuring the introduction of randomness, allowing for optimization by GRPO.
 3.2 Sliding Window as Optimization Scheduler
@@
 In fact, the interval $S$ can be non-fixed during the training process.
-In this section, we will introduce the sliding window to describe the movement of the interval $S$, which leads to a significant improvement in the quality of the generated images.
-Along the discrete denosing timesteps $\{0,1,\ldots,T-1\}$, MixGRPO defines a sliding window $W(l)$ and optimization is only employed at the timesteps within this window.
-$W(l)=\{t_{l},t_{l+1},\dots,t_{l+w-1}\},\quad l\leq T-w$ (10)
+In this section, we will introduce the sliding window to describe the movement of $S$, which leads to a significant improvement in the quality of the generated images.
+Along the denosing time-steps $\{0,1,\ldots,T-1\}$, MixGRPO defines a SDE sliding window $W(l)$ and optimization is only employed at the steps within $W(l)$.
+$W(l)=\{t_{l},t_{l+1},\dots,t_{l+w-1}\},\quad l\leq T-w,$ (10)
 where $l$ is the left boundary of the sliding window, and $w$ is a hyperparameter that represents the window size.
@@
 When the total sampling steps $T=25$, the best performance is achieved with $w=4$, $\tau=25$ and $s=1$.
-The detailed sliding window strategy and the MixGRPO algorithm can be referenced in Algorithm 1
+The detailed sliding window strategy and MixGRPO algorithm can be found in Algorithm 1.
 Restricting the use of SDE sampling within the sliding window not only ensures the diversity of the generated images but also allows the model to concentrate on optimizing the flow within that window.
 Moving along the denoising direction represents the stochasticity of the probability flow from strong to weak, as illustrated in Figure 2.
-This is essentially a greedy strategy and is similar to the RL that assigns discount factors to process rewards (Pitis, 2019; Amit et al., 2020; Hu et al., 2022), which gives greater importance to rewards derived from a larger search space in the earlier process.
+This is essentially a greedy strategy and is similar to the RL that assigns discount factors to process rewards (Pitis, 2019; Amit et al., 2020; Hu et al., 2022b), which gives greater importance to rewards derived from a larger search space in the earlier process.
 The experimental results of different movement strategies in Table 5 demonstrate the validity of this intuition.
@@
 The exponential function $\exp(x)$ calculates $e^{x}$, while the Rectified Linear Unit $\text{ReLU}(x)$ is defined as $\max(0,x)$.
-Table 4 shows that the exponential decay strategy can achieve better results in terms of Pick Score (Kirstain et al., 2023) and ImageReward (Xu et al., 2023), likely because the model focuses on the earlier stages of denoising, which can lead to more significant high-level changes.
+Table 5 shows that the exponential decay strategy can achieve better results in terms of Pick Score (Kirstain et al., 2023) and ImageReward (Xu et al., 2023).
+This may be because the model focuses on the earlier stages of denoising, which can lead to more significant high-level changes, precisely what the human preference alignment reward model emphasizes.
 3.3 Trade-off Between Overhead and Performance
@@
 Therefore, the higher-order ODE solvers e.g., DPM-Solver Series (Lu et al., 2022a; b; Zheng et al., 2023), UniPC (Zhao et al., 2023) designed for DPM sampling acceleration are also applicable to FM.
-We have reformulated DPM-Solver++ (Lu et al., 2022b) to apply it in the FM framework for ODE sampling acceleration and released detailed derivations in Appendix B.
+We have reformulated DPM-Solver++ (Lu et al., 2022b) to apply it in the FM framework for ODE sampling acceleration and released detailed derivations in Appendix C.
 By applying higher-order solvers, we achieve acceleration in the sampling of $\pi_{\theta_{\text{old}}}$ during GRPO training, which is essentially a balance between overhead and performance.
 Excessive acceleration leads to fewer timesteps, which inevitably results in a decline in image generation quality, thereby accumulating errors in the computation of rewards.
-We have found in practice that the 2nd-order DPM-Solver++ is sufficient to provide significant acceleration while ensuring that the generated images align well with human preferences.
-Ultimately, we adopted both progressive and frozen sliding window strategies, proposing MixGRPO-Flash and MixGRPO-Flash*.
-A detailed description of the algorithm can be found in Appendix C.
-These approaches achieve a greater degree of acceleration compared to MixGRPO, while also outperforming DanceGRPO in terms of performance.
+We have found in practice that the 2nd-order DPM-Solver++ is sufficient to provide significant acceleration while ensuring that the generated images align well with human preferences in Table 9.
+Ultimately, we introduced DPM-Solver++ and adopted both progressive and frozen sliding window strategies, proposing MixGRPO-Flash and MixGRPO-Flash*.
+A detailed description of the algorithm can be found in Appendix E.
+These approaches achieve a greater degree of acceleration compared to MixGRPO, while also outperforming DanceGRPO in terms of human preference alignment performance.
 Table 1: Comparison results for overhead and performance.
@@
 The Frozen strategy means that optimization is only employed at the initial denoising steps.
-Method $\text{NFE}_{\pi_{\theta_{\text{old}}}}$ $\text{NFE}_{\pi_{\theta}}$ Iteration Time (s)$\downarrow$ Human Preference Alignment
+[…]
 […]
@@
 Additionally, the average training time per GRPO iteration provides a more accurate reflection of the acceleration effect.
-Performance Evaluation  We used four reward models, namely HPS-v2.1 (Wu et al., 2023), Pick Score (Kirstain et al., 2023), ImageReward (Xu et al., 2023) and Unified Reward Wang et al.
+Performance Evaluation  We used four multiple reward models in conjunction for GRPO, namely HPS-v2.1 (Wu et al., 2023), Pick Score (Kirstain et al., 2023), ImageReward (Xu et al., 2023) and Unified Reward Wang et al.
 (2025), both as reward guidance during training and as evaluation metrics.
@@
 To validate the robustness of MixGRPO, we also followed DanceGRPO and conducted additional comparisons using HPS-v2.1 as a single reward, and combining HPS-v2.1 (Wu et al., 2023) and CLIP Score (Radford et al., 2021) as multi-rewards.
-Figure 4: Qualitative comparison with different training-time sampling steps.
-The performance of MixGRPO does not significantly decrease with the reduction in overhead.
-The Frozen strategy means that optimization is only employed at the initial denoising steps.
 4.2 Implementation Details
-For training-time sampling, we first perform a shift $\tilde{t}=\frac{t}{1-(\tilde{s}-1)t}$ on the uniformly distributed $t=[0,\ldots,T-1]$ and then define $\sigma_{t}=\eta\sqrt{\frac{\tilde{t}}{1-\tilde{t}}}$.
+For training-time sampling, we first perform a shift $\tilde{t}=\frac{t}{1-(\tilde{s}-1)t}$ on the uniformly distributed $t_{i}=\frac{i}{T}$ where $i=[0,\ldots,T-1]$, and then define $\sigma_{t}=\eta\sqrt{\frac{\tilde{t}}{1-\tilde{t}}}$.
 We set $\tilde{s}=3$ and $\eta=0.7$ as scale.
 We set $T=25$ as the total sampling steps.
-For GRPO, the model generates 12 images for each prompt and clips the advantage to the range [5,5][-5,5][ - 5 , 5 ].
+For GRPO, the model generates 12 images for each prompt and clips the advantage to the range [5,5][-5,5].
 It is important to note that we use gradient accumulation over 3 steps, which means that during a single training iteration, there are $\frac{12}{3}=4$ gradient updates.
@@
 Additionally, MixGRPO-Flash enables accelerated sampling of $\pi_{\theta_{\text{old}}}$, and as the overhead decreases, the quality of the generated images still maintains a strong alignment with human preferences.
-Following DanceGRPO (Xue et al., 2025), we also trained and evaluated the model using a single reward model and two reward models, but on the HPDv2 dataset (Wu et al., 2023).
+Following DanceGRPO (Xue et al., 2025), we also trained and evaluated the model using a single reward model, e.g., HPSv2.1 and two reward models, e.g., HPSv2.1 and CLIP Score on the HPDv2 dataset (Wu et al., 2023).
 The results (See Table 3) demonstrate that MixGRPO achieves the best performance on both in-domain and out-of-domain rewards, whether using a single or multiple reward models.
-The visualized results are displayed in Appendix D.
-Table 2: Comparison results for in-domain and out-of-domain reward metrics.
-The results demonstrate that MixGRPO achieves the best performance on both in-domain and out-of-domain rewards, whether using a single or multiple reward models.
+The visualized results are displayed in Figure 6 and Figure 7 of Appendix F.
+To demonstrate the robustness of the MixGRPO, we followed Flow-GRPO (Liu et al., 2025a) and applied the Low-Rank Adaptation (LoRA) (Hu et al., 2022a) method on the Stable Diffusion 3.5 (SD3.5) (Esser et al., 2024).
+We used HPS-v2.1 (Wu et al., 2023), Pick Score (Kirstain et al., 2023), and ImageReward (Xu et al., 2023) as multi-rewards, comparing Flow-DPO (offline/online) (Liu et al., 2025b), and Flow-GRPO (Liu et al., 2025a).
+The results in Table 3 showed that MixGRPO is well-suited for the LoRA, significantly reducing training overhead while outperforming both Flow-DPO and Flow-GRPO in terms of speed and performance.
+The visualization results can be found in Figure 8 of Appendix F.
+Table 2: Comparison results demonstrate that MixGRPO achieves the best performance on both in-domain and out-of-domain rewards.
 Reward Model Method In Domain Out-of-Domain
@@
 MixGRPO 0.349 0.415 0.229 1.416 3.430
+Table 3: Comparison results demonstrate that MixGRPO outperforms Flow-DPO and Flow-GRPO in training efficiency and performance.
+Model RL Method $\text{NFE}_{\pi_{\theta_{\text{old}}}}$ $\text{NFE}_{\pi_{\theta}}$ HPSv2.1 Pick Score ImageReward
+SD3.5-M / / / 0.3066 0.2266 1.1630
+SD3.5-M+DPO LoRA Offline DPO 40 40 0.3043 0.2220 1.4524
+Online DPO 40 40 0.3132 0.2210 1.5001
+SD3.5-M+GRPO LoRA Flow-GRPO 10 10 0.3312 0.2318 1.4572
+MixGRPO 10 4 0.3416 0.2360 1.4854
 4.4 Ablation Experiments
@@
 5 Conclusion
-Although GRPO (Shao et al., 2024) has seen significant success in the language modality, it is still in the early stages of progress in vision (Tong et al., 2025; Xue et al., 2025; Liu et al., 2025).
-Existing flow-matching based GRPO faces challenges such as low sampling efficiency and slow training.
+Although GRPO (Shao et al., 2024) has seen significant success in the language modality, it is still in the early stages of progress in vision (Tong et al., 2025; Xue et al., 2025; Liu et al., 2025a).
+Existing flow-based GRPO faces challenges such as low sampling efficiency and slow training.
 To address these issues, we proposed MixGRPO, a novel training framework that combines SDE and ODE sampling.
 This hybrid approach allows for focused optimization on the SDE sampling flow component, reducing complexity while ensuring accurate reward computation.
-Inspired by the decay factor in reinforcement learning (Hu et al., 2022), we introduce a sliding window strategy for scheduling the optimized denoising steps.
+Inspired by the decay factor in reinforcement learning (Hu et al., 2022b), we introduce a sliding window strategy for scheduling the optimized denoising steps.
 Experimental results confirm the effectiveness of our approach in both single-reward and multi-reward settings.
@@
 We hope MixGRPO will inspire deeper research into post-training for image generation, contributing to the advancement of Artificial General Intelligence (AGI).
-Appendix A Hybrid Inference for Solving Reward Hacking
+6 Acknowledgement
+The authors thank Yiming Cheng (M.Sc., Tsinghua University, 13306339695@163.com) for his valuable feedback on the theoretical derivations and mathematical expressions.
+Appendix A Proof of convergence for Mixed ODE-SDE sampling
+To prove that the mixed ODE-SDE sampling method in Eq.
+(4) has the same convergence as Eq.
+(2), which uses only ODE sampling, referencing Song et al.
+(2020b), we approach this from the perspective of distribution evolution, where the distribution at each time step, e.g., $\frac{\partial{q_{t}(\mathbf{x})}}{\partial{t}}$ must be the same.
+Let the interval for SDE be defined as $S=\left[t_{l},t_{r}\right)\in\left[0,1\right)$.
+Along the denoising direction, when the same initial Gaussian noise distribution $q_{0}(\mathbf{x}_{0})$ is given, the probability distribution evolution in the ODE interval preceding the SDE is completely identical.
+The key point is whether the distribution evolution of the SDE within the interval $S$ is completely equivalent to that of the ODE.
+If they are equivalent, then the ODE interval following the SDE will naturally be equivalent to using only ODE sampling.
+Next, we will provide a detailed proof for this key point.
+Consider the SDE Eq. (3) in the interval $S$, which possesses the following form:
+$\mathrm{d}\mathbf{x}=[f(\mathbf{x},t)-g^{2}(t)\nabla_{\mathbf{x}}\log{q_{t}(\mathbf{x})}]\mathrm{d}t+g(t)\mathrm{d}\mathbf{w},\quad t\in S.$ (12)
+The marginal probability density $q_{t}(\mathbf{x}_{t})$ evolves according to Kolmogorov’s equation (Fokker-Planck equation) (Øksendal, 2003)
+$\frac{\partial q_{t}(\mathbf{x})}{\partial t}=-\nabla_{\mathbf{x}}\cdot\big[\big(f(\mathbf{x},t)-g^{2}(t)\nabla_{\mathbf{x}}\log q_{t}(\mathbf{x})\big)q_{t}(\mathbf{x})\big]+\frac{1}{2}g^{2}(t)\nabla^{2}_{\mathbf{x}}q_{t}(\mathbf{x})$ (13)
+[…]
+$\frac{\partial q_{t}(\mathbf{x})}{\partial t}$ $=-\nabla_{\mathbf{x}}\cdot\big[f(\mathbf{x},t)q_{t}(\mathbf{x})-g^{2}(t)\nabla_{\mathbf{x}}q_{t}(\mathbf{x})\big]+\frac{1}{2}g^{2}(t)\nabla^{2}_{\mathbf{x}}q_{t}(\mathbf{x})$
+$=-\nabla_{\mathbf{x}}\cdot[f(\mathbf{x},t)q_{t}(\mathbf{x})-\frac{1}{2}g^{2}(t)\nabla_{\mathbf{x}}q_{t}(\mathbf{x})]$
+$=-\nabla_{\mathbf{x}}\cdot\big[\underbrace{\big(f(\mathbf{x},t)-\frac{1}{2}g^{2}(t)\nabla_{\mathbf{x}}\log q_{t}(\mathbf{x})\big)}_{f_{\text{ODE}}(\mathbf{x},t)}q_{t}(\mathbf{x})\big].$ (14)
+The Eq. (A) is indeed the Fokker-Planck equation of the ODE Eq. (2). Therefore, within the interval $S$, the distribution evolution of SDE and ODE sampling is consistent.
+Appendix B Hybrid Inference for Solving Reward Hacking
 To address the reward hacking issue, we employ hybrid inference to sample with both the raw and post-trained models, and introduce the hybrid percent $p_{\text{mix}}$.
@@
 Figure 4: Qualitative comparison with different hybrid inference percentages
-Appendix B DPM-Solver++ for Recitified Flow
-The DPM-Solver++ algorithm (Lu et al., 2022b) is originally designed for the $\mathbf{x}_{0}\text{-}prediction$ diffusion model (Rombach et al., 2022), where the model outputs the denoised feature $\mathbf{x}_{0}$ based on the noisy feature $\mathbf{x}_{t}$, the time condition $t$, and the text condition $c$.
+Appendix C DPM-Solver++ for Recitified Flow
+For clarity and to avoid ambiguity between continuous time and discrete steps, we adopt the following notation in this section.
+We denote the discrete time steps by an index $i\in\{0,1,\dots,T-1\}$, where $T$ is the total number of sampling steps.
+The continuous time corresponding to step $i$ is denoted by $t_{i}=\frac{i}{T}\in[0,1)$.
+The DPM-Solver++ algorithm (Lu et al., 2022b) is originally designed for the $\mathbf{x}_{0}\text{-}prediction$ diffusion model (Rombach et al., 2022), where the model outputs the denoised feature $\mathbf{x}_{0}$ based on the noisy feature $\mathbf{x}_{t_{i}}$, the time condition $t_{i}$ and the text condition $c$.
 According to the definition of Rectified Flow (RF) (Liu et al., 2022), there is the following transfer equation:
-$\mathbf{x}_{t}=t\mathbf{x}_{1}+(1-t)\mathbf{x}_{0}$ (12)
-According to the theory of stochastic interpolation (Albergo et al., 2023), RF effectively approximates $\mathbf{x}_{1}-\mathbf{x}_{0}$ by modeling $\mathbf{v}_{t}$:
-$\mathbf{v}_{t}=\mathbf{x}_{1}-\mathbf{x}_{0}$ (13)
+$\mathbf{x}_{t_{i}}=t_{i}\mathbf{x}_{1}+(1-t_{i})\mathbf{x}_{0}.$ (15)
+According to the theory of stochastic interpolation (Albergo et al., 2023), RF effectively approximates $\mathbf{x}_{1}-\mathbf{x}_{0}$ by modeling $\mathbf{v}_{t_{i}}$:
+$\mathbf{v}_{t_{i}}=\mathbf{x}_{1}-\mathbf{x}_{0}.$ (16)
 Based on Eq. (15) and Eq. (16), we obtain the following relationship:
-$\mathbf{x}_{0}=\mathbf{x}_{t}-\mathbf{v}_{t}t$ (14)
+$\mathbf{x}_{0}=\mathbf{x}_{t_{i}}-\mathbf{v}_{t_{i}}t_{i}.$ (17)
 By using a neural network for approximation, we establish the relationship between RF and the $\mathbf{x}_{0}\text{-}prediction$ model:
-$\mathbf{x}_{\theta}=\mathbf{x}_{t}-\mathbf{v}_{\theta}t$ (15)
-Taking the multistep second-order solver as an example (see Algorithm 2 in  (Lu et al., 2022b)), we derive the corrected $\mathbf{x}_{\theta}$ for the RF sampling process as $\mathbf{D}_{t}$:
-$\mathbf{D}_{t}\leftarrow\left(1+\frac{h_{t}}{2h_{t-1}}\right)\left(\mathbf{x}_{t-1}-\mathbf{v}_{\theta}(\mathbf{x}_{t-1},t-1,c)\frac{t-1}{T}\right)$ (16)
-$-\frac{h_{t}}{2h_{t-1}}\left(\mathbf{x}_{t-2}-\mathbf{v}_{\theta}(\mathbf{x}_{t-2},t-2,c)\frac{t-2}{T}\right)$
-where $T$ is the total sampling steps and $h_{t}=\lambda_{t}-\lambda_{t-1}$.
-$\lambda_{t}$ is the log-signal-to-noise-ratio (log-SNR) and is defined in RF as $\lambda_{t}:=\log\left(\frac{1-t}{t}\right)$.
+$\mathbf{x}_{\theta}(\mathbf{x}_{i},t_{i},c)=\mathbf{x}_{i}-\mathbf{v}_{\theta}(\mathbf{x}_{i},t_{i},c)\cdot t_{i}.$ (18)
+Taking the multistep second-order DPMSolver++ as an example (see Algorithm 2 in  (Lu et al., 2022b)), we derive the corrected $\mathbf{x}_{\theta}$ for the RF sampling process as $\mathbf{D}_{i}$:
+$\mathbf{D}_{i}\leftarrow\left(1+\frac{h_{i}}{2h_{i-1}}\right)\left(\mathbf{x}_{i-1}-\mathbf{v}_{\theta}(\mathbf{x}_{i-1},t_{i-1},c)\cdot t_{i-1}\right)$ (19)
+$\quad-\frac{h_{i}}{2h_{i-1}}\left(\mathbf{x}_{i-2}-\mathbf{v}_{\theta}(\mathbf{x}_{i-2},t_{i-2},c)\cdot t_{i-2}\right),$
+where $h_{i}=\lambda_{t_{i}}-\lambda_{t_{i-1}}$.
+The continuous time $t_{i}$ corresponds to the discrete step $i$ over a total of $T$ sampling steps.
+The term $\lambda_{t_{i}}$ is the log-signal-to-noise-ratio (log-SNR) and is defined in RF as:
+$\lambda_{t_{i}}:=\log\left(\frac{1-t_{i}}{t_{i}}\right).$ (20)
 Based on the exact discretization formula for the probability flow ODE proposed in DPM-Solver++ (Eq. (9) in (Lu et al., 2022b)), we can derive the final transfer equation:
-$\mathbf{x}_{t}\leftarrow\frac{t}{t-1}\mathbf{x}_{t-1}-(1-t)\left(e^{-h_{t}}-1\right)\mathbf{D}_{t},\quad 1\leq t<T$ (17)
-Appendix C MixGRPO-Flash Algorithm
-MixGRPO-Flash accelerates the ODE sampling that does not contribute to the calculation of the policy ratio after the sliding window by using DPM-Solver++ in the Eq.
-(17).
-We introduce a compression rate $\tilde{r}$ such that the ODE sampling after the window only requires $(T-l-w)\tilde{r}$ time steps.
-And the total time-steps is $\tilde{T}=l+w+(T-l-w)\tilde{r}$ The final algorithm is as follows:
-Algorithm 2 MixGRPO-Flash Training Process
-1:initial policy model $\pi_{\theta}$; reward models $\{R_{k}\}_{k=1}^{K}$; prompt dataset $\mathcal{C}$; total sampling steps $\tilde{T}$; number of samples per prompt $N$; ODE compression rate $\tilde{r}$
+$\mathbf{x}_{i}\leftarrow\frac{t_{i}}{t_{i-1}}\mathbf{x}_{i-1}-(1-t_{i})\left(e^{-h_{i}}-1\right)\mathbf{D}_{i},\quad 1\leq i<T.$ (21)
+Appendix D MixGRPO Algorithm
+Taking the progressive movement strategy of the sliding window as an example, the MixGRPO Algorithm 1 is as follows.
+MixGRPO first sets the left boundary $l$ of the sliding window $W(l)$ at $t=0$ of the reverse process, which corresponds to the first sampling time-step.
+A GRPO iteration consists of two phases: sampling and optimization.
+In the sampling phase, mixed ODE-SDE sampling is performed for the $N$ samples within the group, using SDE within the window and ODE outside.
+Then, the reward and advantage for the $N$ samples are calculated.
+For the multi-reward strategy, the advantage for each reward model is computed separately and then aggregated to obtain the final advantage.
+In the optimization phase, the GRPO loss (Shao et al., 2024) is computed and optimized only at the sampling timesteps within the SDE window $W(l)$.
+After each GRPO iteration, the current training steps are checked against the conditions for moving the window, determining whether to shift the window by the window stride $s$.
+Algorithm 1 MixGRPO Training Process
+1:initial policy model $\pi_{\theta}$; reward models $\{R_{k}\}_{k=1}^{K}$; prompt dataset $\mathcal{C}$; total sampling steps $T$; number of samples per prompt NN;
 2:sliding window $W(l)$, window size $w$, shift interval $\tau$, window stride $s$
@@
 8:   Init the same noise $\mathbf{x_{0}}\sim\mathcal{N}(0,\mathbf{I})$
-9:   for generate $i$-th image from $i=1$ todo
+9:   for generate $i$-th image from $i=1$ to N do
+10:     for sampling timestep $t=0$ to T1T-1 do $\triangleright$ $\pi_{\theta_{\text{old}}}$ mixed sampling loop
+11:      if $t\in W(l)$ then
+12:        Use SDE Sampling to get $\mathbf{x}^{i}_{t+1}$
+13:      else
+14:        Use ODE Sampling to get $\mathbf{x}^{i}_{t+1}$
+15:      end if
+16:     end for
+17:   end for
+18:   for $i$-th image from $i=1$ to N do
+19:     Calculate multi-reward advantage: $A_{i}\leftarrow\sum_{k=1}^{K}\frac{R(\mathbf{x}^{i}_{T},\mathbf{c})^{i}_{k}-\mu_{k}}{\sigma_{k}}$
+20:   end for
+21:   for optimization timestep $t\in W(l)$ do $\triangleright$ optimize policy model $\pi_{\theta}$
+22:     Update policy model via gradient ascent: $\theta\leftarrow\theta+\eta\nabla_{\theta}\mathcal{J}$
+23:   end for
+24:  end for
+25:  if $m\bmod\tau~\text{is}~0$ then $\triangleright$ move sliding window
+26:   $l\leftarrow\min(l+s,~T-w)$
+27:  end if
+28:end for
+Appendix E MixGRPO-Flash Algorithm
+MixGRPO-Flash Algorithm 2 accelerates the ODE sampling that does not contribute to the calculation of the policy ratio after the sliding window by using DPM-Solver++ in the Eq.
+(21).
+We introduce a compression rate $\tilde{r}$ such that the ODE sampling after the window only requires $(T-l-w)\tilde{r}$ time steps.
+And the total time-steps is $\tilde{T}=l+w+(T-l-w)\tilde{r}$ The final algorithm is as follows:
+Note that when using MixGRPO-Flash*, the frozen strategy is applied, with the left boundary of the sliding window $l\equiv 0$.
+The theoretical speedup of the training-time sampling can be described as follows:
+$S=\frac{T}{w+(T-w)\tilde{r}}$. (22)
+For MixGRPO-Flash, since the sliding window moves according to the progressive strategy during training, the average speedup can be expressed in the following form:
+$S=\frac{T}{\mathbb{E}_{l}\left(w+l+\lceil(T-w-l)\tilde{r}\rceil\right)}<\frac{T}{w+(T-w)\tilde{r}}$. (23)
+Algorithm 2 MixGRPO-Flash Training Process
+1:initial policy model $\pi_{\theta}$; reward models $\{R_{k}\}_{k=1}^{K}$; prompt dataset $\mathcal{C}$; total sampling steps $\tilde{T}$; number of samples per prompt NN; ODE compression rate $\tilde{r}$
+2:sliding window $W(l)$, window size $w$, shift interval $\tau$, window stride $s$
+3:Init left boundary of $W(l)$: $l\leftarrow 0$
+4:for training iteration $m=1$ to $M$ do
+5:  Sample batch prompts $\mathcal{C}_{b}\sim\mathcal{C}$
+6:  Update old policy model: $\pi_{\theta_{\text{old}}}\leftarrow\pi_{\theta}$
+7:  for each prompt $\mathbf{c}\in\mathcal{C}_{b}$ do
+8:   Init the same noise $\mathbf{x_{0}}\sim\mathcal{N}(0,\mathbf{I})$
+9:   for generate $i$-th image from $i=1$ to N do
 10:     for sampling timestep $t=0$ to $\tilde{T}-1$ do
@@
 14:        Use SDE sampling to get $\mathbf{x}^{i}_{t+1}$
-15:      $\triangleright$ DPM-Solver++
+[…]
 16:        Use higher-order ODE sampling to get $\mathbf{x}^{i}_{t+1}$
@@
 18:     end for
-19:     Calculate multi-reward advantage: $A_{i}\leftarrow\sum_{k=1}^{K}\frac{R(\mathbf{x}^{i}_{\tilde{T}},\mathbf{c})^{i}_{k}-\mu_{k}}{\sigma_{k}}$
-20:   end for
-21:   for optimization timestep $t\in W(l)$ do $\triangleright$ optimize policy model $\pi_{\theta}$
-22:     Update policy model via gradient ascent: $\theta\leftarrow\theta+\eta\nabla_{\theta}\mathcal{J}$
-23:   end for
-24:  end for
-25:  if use MixGRPO-Flash* then $\triangleright$ move sliding window
-26:   $l\leftarrow 0$
-27:  else
-28:   if $m\bmod\tau~\text{is}~0$ then
-29:     $l\leftarrow\min(l+s,~T-w)$
-30:   end if
-31:  end if
-32:end for
-Note that when using MixGRPO-Flash*, the frozen strategy is applied, with the left boundary of the sliding window $l\equiv 0$.
-The theoretical speedup of the training-time sampling can be described as follows:
-$S=\frac{T}{w+(T-w)\tilde{r}}$ (18)
-For MixGRPO-Flash, since the sliding window moves according to the progressive strategy during training, the average speedup can be expressed in the following form:
-$S=\frac{T}{\mathbb{E}_{l}\left(w+l+\lceil(T-w-l)\tilde{r}\rceil\right)}<\frac{T}{w+(T-w)\tilde{r}}$ (19)
-Appendix D More visualized results
+19:   end for
+20:   for $i$-th image from $i=1$ to N do
+21:     Calculate multi-reward advantage: $A_{i}\leftarrow\sum_{k=1}^{K}\frac{R(\mathbf{x}^{i}_{\tilde{T}},\mathbf{c})^{i}_{k}-\mu_{k}}{\sigma_{k}}$
+22:   end for
+23:   for optimization timestep $t\in W(l)$ do $\triangleright$ optimize policy model $\pi_{\theta}$
+24:     Update policy model via gradient ascent: $\theta\leftarrow\theta+\eta\nabla_{\theta}\mathcal{J}$
+25:   end for
+26:  end for
+27:  if use MixGRPO-Flash* then $\triangleright$ move sliding window
+28:   $l\leftarrow 0$
+29:  else
+30:   if $m\bmod\tau~\text{is}~0$ then
+31:     $l\leftarrow\min(l+s,~T-w)$
+32:   end if
+33:  end if
+34:end for
+Appendix F More visualized results
+Figure 5: Qualitative comparison with different training-time sampling steps.
+The performance of MixGRPO does not significantly decrease with the reduction in overhead.
+The Frozen strategy means that optimization is only employed at the initial denoising steps.
 Figure 6: Comparison of the visualization results of FLUX, DanceGRPO, and MixGRPO under HPS-v2.1 as the reward model.
-Figure 7: Comparison of the visualization results of FLUX, DanceGRPO, and MixGRPO under HPS-v2.1 and CLIP Score as the reward models.
+Figure 7: Comparison of the visualization results of FLUX, DanceGRPO, and MixGRPO under HPS-v2.1 and CLIP Score as multi-reward models.
+Figure 8: Comparison of the visualization results of SD3.5-M, offline DPO, online DPO, Flow-GRPO and MixGRPO under HPS-v2.1, Pick Score and ImageReward as multi-reward models.

### v3

@@
 Abstract
-Although GRPO substantially enhances flow matching models in human preference alignment of image generation, methods such as DanceGRPO still exhibit inefficiency due to the necessity of sampling and optimizing over all denoising steps specified by the Markov Decision Process (MDP).
+Although GRPO substantially enhances flow matching models in human preference alignment of image generation, methods such as FlowGRPO and DanceGRPO still exhibit inefficiency due to the necessity of sampling and optimizing over all denoising steps specified by the Markov Decision Process (MDP).
 In this paper, we propose MixGRPO, a novel framework that leverages the flexibility of mixed sampling strategies through the integration of stochastic differential equations (SDE) and ordinary differential equations (ODE).
@@
 This design confines sampling randomness to the time-steps within the window, thereby reducing the optimization overhead, and allowing for more focused gradient updates to accelerate convergence.
-Additionally, as time-steps beyond the sliding window are not involved in optimization, higher-order solvers are supported for sampling.
+Additionally, as time-steps beyond the sliding window are not involved in optimization, higher-order solvers are supported for faster sampling.
 So we present a faster variant, termed MixGRPO-Flash, which further improves training efficiency while achieving comparable performance.
@@
 Notably, MixGRPO-Flash further reduces training time by 71%.
-Codes and models are available at MixGRPO.
+[^1] Code is available in https://github.com/Tencent-Hunyuan/MixGRPO.
 1 Introduction
-Recent advances (Liu et al., 2025a; Xue et al., 2025; Liu et al., 2025b; Li et al., 2025; Xu et al., 2023) in Text-to-Image (T2I) tasks have demonstrated that probability flow models can achieve improved performance by incorporating Reinforcement Learning from Human Feedback (RLHF) (Ouyang et al., 2022) strategies during the post-training stage to maximize rewards.
-Specifically, methods (Liu et al., 2025a; Xue et al., 2025) based on Group Relative Policy Optimization (GRPO) (Shao et al., 2024), have recently been studied, achieving optimal alignment with human preferences.
-Current GRPO methods in probability flow models, e.g., Flow-GRPO (Liu et al., 2025a), DanceGRPO (Xue et al., 2025), leverage Stochastic Differential Equations (SDE) sampling at every denoising step to introduce randomness, addressing reliance on stochastic exploration in RLHF.
-They model the entire denoising process as a Markov Decision Process (MDP) in a stochastic environment, using GRPO to optimize the state-action sequence.
-However, the need to optimize all denoising steps not only increases overhead but also leads to inconsistent gradient descent, resulting in inefficient training.
-Specifically, to compute the policy ratio, it is essential to perform full-step sampling independently with the old policy model $\pi_{\theta_{\text{old}}}$ and new one $\pi_{\theta}$.
-While DanceGRPO proposes to randomly select a subset of steps at a fixed ratio to optimize, our empirical analysis in Figure 1 demonstrates a substantial performance degradation as the subset size is reduced.
-Figure 1: Performance comparison for different numbers of denoising steps optimized.
-The performance improvement of DanceGRPO relies on more steps optimized.
-MixGRPO achieves optimal performance while requiring only 4 steps.
-To address these issues, we propose MixGRPO, which achieves more efficient stochastic exploration while enabling optimization over fewer denoising steps.
-Specifically, we employ a mixed ODE-SDE strategy, applying SDE sampling to a denoising sub-interval and Ordinary Differential Equations (ODE) sampling to the rest, confining randomness to the SDE interval.
-In this way, fewer time-steps are needed for GRPO optimization, without compromising image quality for reward computation.
-Besides, we introduce a sliding window strategy for the SDE interval that moves along the denoising steps as the training process progresses.
-Unlike random selection, this scheduling strategy orders optimization from high to low denoising levels, which aligns with the intuition of applying temporal discount factors to rewards in Reinforcement Learning (RL) (Pitis, 2019; Amit et al., 2020; Hu et al., 2022b).
-MixGRPO prioritizes optimizing the initial timesteps, which involve the most significant noise removal and entail a larger exploration space (see Figure 2).
-As training progresses, MixGRPO continuously moves the SDE window to narrow the exploration space and achieve global optimization.
-Finally, we find that higher-order ODE solvers, e.g., DPMSolver++ (Lu et al., 2022b), can significantly accelerate training-time sampling, with negligible performance degradation, as there is no need for the posterior probability distribution after the sliding window.
-We trained and evaluated MixGRPO by using HPS-v2.1 (Wu et al., 2023), Pick Score (Kirstain et al., 2023), ImageReward (Xu et al., 2023), and Unified Reward (Wang et al., 2025) as reward models (RMs) and metrics.
-We also quantified the overhead in terms of the number of function evaluations (NFE) and time consumption during training.
-During the training process, we fine-tuned based on FLUX.1-dev (Labs, 2024) and compared performance using either a single RM or multiple RMs as guidance, assessing the results for both in-domain and out-of-domain metrics.
-Specifically, trained and evaluated on the HPDv2 dataset (Wu et al., 2023), MixGRPO outperforms DanceGRPO across all metrics, particularly improving the ImageReward (Xu et al., 2023) from 1.088 to 1.629, surpassing DanceGRPO’s score of 1.436, while generating images with enhanced semantic quality, aesthetics, and reduced distortion.
+Recent advances [17, 42, 18, 41, 48] in Text-to-Image (T2I) tasks have demonstrated that probability flow models can achieve improved performance by incorporating Reinforcement Learning from Human Feedback (RLHF) [25] strategies during the post-training stage to maximize rewards.
+Specifically, methods [17, 42] based on Group Relative Policy Optimization (GRPO) [32], have recently been studied, achieving optimal alignment with human preferences.
+Current GRPO methods in probability flow models, such as Flow-GRPO and DanceGRPO, combine a rollout phase followed by an online optimization phase.
+During rollouts, they employ Stochastic Differential Equation (SDE) sampling at each denoising step to introduce randomness, thereby enabling the stochastic exploration required by RLHF.
+This allows the entire denoising process to be framed as a Markov Decision Process (MDP) in a stochastic environment, where GRPO is then applied to optimize the complete state-action sequence in the online optimization phase.
+However, the requirement to optimize the entire denoising steps introduces two challenges, prohibitive computational overhead and unfocused and inefficient optimization.
+i) The overhead arises because computing the policy ratio requires separate full-trajectory sampling from both the old ($\pi_{\theta_{\text{old}}}$) and new policies ($\pi_{\theta}$).
+Although DanceGRPO proposes a workaround by optimizing a random subset of steps, our findings in Figure 1 reveal a critical trade-off: reducing the subset size to save computation severely compromises model performance.
+ii) Early-step gradients prioritize global structure while late-step gradients focus on fine details, leading to conflicting update signals.
+This may result in an inefficient training process.
+Figure 1: Comparison of our MixGRPO and DanceGRPO with varying denoising steps to be optimized. MixGRPO achieves higher performance with lower overhead.
+To address these issues, we propose MixGRPO, a method that optimizes a strategic subset of denoising steps to drastically reduce computational overhead while ensuring a more focused and efficient optimization.
+Specifically, we employ a mixed ODE-SDE strategy, applying SDE sampling to a denoising sub-interval and Ordinary Differential Equations (ODE) sampling to the rest.
+Figure 2: (Left) Methods Comparison.
+By employing the mixed ODE-SDE and sliding window strategy, MixGRPO reduces the number of optimized timesteps and achieves more efficient training than DanceGRPO [42].
+(Right) SDE Exploration Analysis.
+As training progresses, MixGRPO gradually slides the SDE window from a high-noise to a low-noise regime, causing the exploration space to shrink and the distribution of sampled images to transition from dispersed to concentrated.
+The policy optimization is confined only to the stochastic SDE sub-interval.
+This interval operates as a sliding window, systematically advancing from low-SNR (signal-to-noise ratio) steps to high-SNR denoising steps as training progresses.
+This curriculum-based approach is conceptually analogous to temporal discounting in RL [26, 2, 10], as it prioritizes the optimization of initial, high-impact steps that establish global structure from a vast exploration space (Figure 2 Right), before progressively shifting focus to the refinement of local details.
+This selective optimization not only reduces the computational burden but also fosters a more dedicated exploration.
+Furthermore, this decoupling allows us to accelerate the deterministic ODE portions with fast solvers (e.g., DPMSolver++ [22]), since their posterior distributions are irrelevant to the optimization, thereby reducing rollout time without compromising final image quality.
+To evaluate the performance and efficiency of MixGRPO, we conduct comprehensive experiments by implementing our method within the frameworks of both Flow-GRPO [17] and DanceGRPO [42].
+We use a set of reward models (HPS-v2.1 [40], Pick Score [12], ImageReward [41], Unified Reward [37], etc) that serve as both training guidance and evaluation metrics.
+The evaluation compares single versus multi-RM guidance and assesses both in-domain and out-of-domain performance.
+Alongside performance, computational overhead is tracked via the number of function evaluations (NFE) and training time.
+Specifically, trained and evaluated on the HPDv2 dataset [40], MixGRPO outperforms DanceGRPO across all metrics, particularly improving the ImageReward [41] from 1.088 to 1.629, surpassing DanceGRPO’s score of 1.436, while generating images with enhanced semantic quality, aesthetics, and reduced distortion.
 Furthermore, MixGRPO reduces the training time of DanceGRPO by nearly 50%.
-In addition, MixGRPO-Flash utilizes DPMSolver++ (Lu et al., 2022b) to accelerate the sampling of $\pi_{\theta_{\text{old}}}$, reducing training time by 71%.
-To summarize, the key contributions of our work are outlined below:
-  • We propose a mixed ODE-SDE GRPO training framework for flow-based models, which alleviates the overhead bottleneck by streamlining the optimization process within the MDP.
-  • We introduce a sliding window strategy to sample the denoising timesteps for model optimization, aligning with the RL intuition of transitioning from harder to easier search spaces, significantly enhancing performance.
-  • Our method enables the use of higher-order ODE solvers to accelerate $\pi_{\theta_{\text{old}}}$ sampling during GRPO training, achieving more significant speed improvements with comparable performance.
-  • Comprehensive experiments were carried out on multiple rewards and the results demonstrate that MixGRPO achieves substantial gains on various evaluation metrics, while significantly reducing training overhead.
-Figure 2: Visualization of t-SNE (Van der Maaten & Hinton, 2008) for images sampled with different strategies.
-Employing SDE sampling in the early stages of the denoising process results in a more discrete data distribution.
+In addition, MixGRPO-Flash utilizes DPMSolver++ [22] to accelerate the sampling of $\pi_{\theta_{\text{old}}}$, reducing training time by 71%.
+To summarize, the key contributions are outlined below:
+  • We propose a mixed ODE-SDE framework for GRPO training with flow-based models.
+This approach alleviates computational overhead by confining the stochastic optimization to a specific sub-interval of the Markov Decision Process.
+  • We introduce an SDE sliding window to dynamically schedule the optimization steps.
+By guiding the optimization from broad exploration to fine-grained refinement, this strategy significantly enhances performance.
+  • Our hybrid framework enables the use of high-order ODE solvers to accelerate the deterministic sampling of $\pi_{\theta_{\text{old}}}$ during GRPO training.
+This yields substantial speedups with negligible performance degradation.
+  • Comprehensive experiments across multiple frameworks demonstrate the versatility of MixGRPO.
+Our method consistently achieves substantial performance gains while significantly reducing training overhead, highlighting its broad applicability and effectiveness.
 2 Related Work
 2.1 RL For Image Generation
-Inspired by Proximal Policy Optimization (PPO) (Schulman et al., 2017), early works (Fan & Lee, 2023; Black et al., 2023; Fan et al., 2023; Lee et al., 2023) integrated reinforcement learning (RL) into diffusion models by optimizing the score function (Song et al., 2020b) through policy gradient methods, thereby enabling the generation of images that better align with human preferences.
-Subsequently,  Wallace et al.
-(2024) introduced offline-Direct Preference Optimization (DPO) to T2I tasks for the first time.
+Inspired by Proximal Policy Optimization (PPO) [31], early works [5, 3, 6, 14] integrated reinforcement learning (RL) into diffusion models by optimizing the score function [34] through policy gradient methods, thereby enabling the generation of images that better align with human preferences.
+Subsequently,  [36] introduced offline-Direct Preference Optimization (DPO) to T2I tasks for the first time.
 This allows diffusion models to directly learn from human feedback and validates its effectiveness on large-scale models.
-Due to the tendency of offline win-lose pair data to shift the model away from its original distribution, some works (Yuan et al., 2024; Liang et al., 2025) have adopted online methods, continuously adjusting sampling trajectories through step-aware preference models during training to achieve improved performance.
-Recently, GRPO-based works e.g.,  Tong et al.
-(2025), Flow-GRPO (Liu et al., 2025a) and DanceGRPO (Xue et al., 2025), have elevated RL-enhanced image generation to new heights.
-Specifically, Flow-GRPO (Liu et al., 2025a) and DanceGRPO (Xue et al., 2025) introduced GRPO to flow matching models, allowing divergent sampling by transforming the ODE into an equivalent SDE.
-They also identified the overhead caused by full-step sampling within a group as a bottleneck and sought to address it by reducing or randomly selecting denoising steps.
-However, these approaches do not fundamentally address the issue.
+Due to the tendency of offline win-lose pair data to shift the model away from its original distribution, some works [45, 15] have adopted online methods, continuously adjusting sampling trajectories through step-aware preference models during training to achieve improved performance.
+Recently, GRPO-based works e.g.,  [35], Flow-GRPO [17] and DanceGRPO [42], have elevated RL-enhanced image generation to new heights.
+Specifically, they introduced GRPO to flow matching models, allowing divergent sampling by transforming the ODE into an equivalent SDE.
+DanceGRPO [42] also identified the overhead caused by full-step optimization as a bottleneck and sought to address it by reducing or randomly selecting a subset.
+However, this approach does not fundamentally address the issue.
 We hope to delve into the essence of GRPO on the probability flow and provide deeper insights through mixed sampling techniques and optimization scheduling.
 2.2 Sampling Methods for Probability Flow
-DDPM (Ho et al., 2020) first proposed training a probabilistic model to reverse each step of noise corruption and utilized probability flow SDE for sampling, enabling the generation of realistic images.
-However, this often requires thousands of steps, resulting in significant overhead.
-DDIM (Song et al., 2020a) introduced deterministic sampling and proposed a probability ODE sampling approach, reducing the number of sampling steps to around 100.
-Subsequently, inspired by the Fokker-Planck equation (Risken & Risken, 1996), Song et al.
-(2020b) established a unification of SDE and ODE sampling methods from the perspective of the score function.
-Then, more higher-order ODE solvers were proposed e.g., DPM-Solver (Lu et al., 2022a) and DPM-Solver++ (Lu et al., 2022b), which utilize multistep methods for differential discretization.
-These approaches significantly reduce the number of sampling steps to around 10 while preserving accuracy.
-Higher-performance solvers (Zheng et al., 2023; Zhao et al., 2023) continue to be proposed; however, the gains are relatively marginal and have ultimately been replaced by the distillation method (Salimans & Ho, 2022; Yin et al., 2024).
-During the same period, flow matching models (Lipman et al., 2022; Esser et al., 2024) simplified and stabilized training by predicting the vector field velocity, enabling deterministic sampling with ODEs under 50 steps.
-Recent theoretical works (Gao et al., 2024; Albergo et al., 2023) has proven that the sampling method of flow matching is equivalent to DDIM, and demonstrated that flow matching models share the same equivalent SDE and ODE formulations as diffusion models.
-This provides important theoretical support and insights for our work, and we may explore interleaved sampling of SDE and ODE in probability flow models as a potential approach.
+The development of generative sampling began with DDPM [8], which used a slow, SDE-based probability flow requiring thousands of steps.
+To accelerate this, DDIM [33] introduced a deterministic ODE-based approach, reducing sampling steps to around 100.
+This SDE / ODE duality was later unified by score-based models [34], paving the way for higher-order ODE solvers like DPM-Solver [21, 22] that reduced steps to 10.
+Higher-performance solvers [47, 46] continue to be proposed; however, the gains are relatively marginal and have ultimately been replaced by the distillation method [30, 44].
+Concurrently, flow matching models [16, 4] simplified training by directly predicting the vector field, also enabling fast ODE sampling.
+Crucially, recent theoretical work [7, 1] has proven that flow matching and diffusion models share equivalent SDE and ODE formulations.
+This unification provides the theoretical foundation for our work, in which we explore an interleaved SDE and ODE sampling strategy within these probability flow models, thereby confining stochasticity to the SDE intervals in RL optimization, which shortens the effective MDP horizon and reduces training overhead.
 3 Method
 3.1 Mixed ODE-SDE Sampling in GRPO
-According to Flow-GRPO (Liu et al., 2025a), the SDE sampling in flow matching can be framed as a Markov Decision Process (MDP) $\left(\mathcal{S},\mathcal{A},\rho_{0},P,\mathcal{R}\right)$ in a stochastic environment.
-The agent produces a trajectory during the discrete sampling process defined as $\Gamma=(\mathbf{s}_{0},\mathbf{a}_{0},\mathbf{s}_{1},\mathbf{a}_{1},\ldots,\mathbf{s}_{T},\mathbf{a}_{T})$, where the reward is provided only at the final step by the reward model, specifically $\mathcal{R}(\mathbf{s}_{i},\mathbf{a}_{i})\overset{\triangle}{=}R(\mathbf{x}_{T},c)$ if $i=T$, and 0 otherwise.
-In MixGRPO, we propose a hybrid sampling method that combines SDE and ODE.
-MixGRPO defines a time interval $S=[t_{l},t_{r})\in[0,1)$, which corresponds to a subinterval of denoising timesteps, such that $0\leq l<r\leq T$ and $t_{i}=\frac{i}{T}$.
-We use SDE sampling within the interval $S$ and ODE sampling outside, while $S$ shifts along the denoising direction throughout the training process (See Figure 2).
-MixGRPO restricts the agent’s stochastic exploration space to the interval $S$, shortening the sequence length of the MDP to a subset $\Gamma_{\text{MixGRPO{}}}=(\mathbf{s}_{l},\mathbf{a}_{l},\mathbf{s}_{l+1},\mathbf{a}_{l+1},\ldots,\mathbf{s}_{r},\mathbf{a}_{r})$ and requires reinforcement learning (RL) optimization only on this subset:
-$\max_{\theta}\mathbb{E}_{\Gamma_{\text{MixGRPO}}\sim\pi_{\theta}}\left[\sum_{t\in S}\big(\mathcal{R}\left(\mathbf{s}_{t},\mathbf{a}_{t}\right)-\beta D_{KL}\left(\pi\left(\cdot|\mathbf{s}_{t}\right)||\pi_{\text{ref}}\left(\cdot|\mathbf{s}_{t}\right)\right)\big)\right],$ (1)
-where $\mathcal{R}(\mathbf{s}_{i},\mathbf{a}_{i})\overset{\triangle}{=}R(\mathbf{x}_{T},c)$ if $i=r$, and 0 otherwise.
-MixGRPO reduces computational overhead while also lowering the difficulty of optimization.
-Next, we derive the specific sampling form and optimization objective of MixGRPO.
-For a deterministic reverse probability flow ODE (Song et al., 2020b), it takes the following form:
-$\frac{\mathrm{d}\mathbf{x}_{t}}{\mathrm{d}t}=f(\mathbf{x}_{t},t)-\frac{1}{2}g^{2}(t)\nabla_{\mathbf{x}_{t}}\log{q_{t}(\mathbf{x}_{t})},\quad\mathbf{x}_{0}\sim q_{0}(\mathbf{x}_{0}),$ (2)
+According to Flow-GRPO [17], the SDE sampling in flow matching can be framed as a Markov Decision Process (MDP) $\left(\mathcal{S},\mathcal{A},\rho_{0},P,\mathcal{R}\right)$.
+The agent produces a trajectory during the discrete sampling process defined as $\Gamma=(\mathbf{s}_{0},\mathbf{a}_{0},\mathbf{s}_{1},\mathbf{a}_{1},\ldots,\mathbf{s}_{T},\mathbf{a}_{T})$, where the reward is provided only at the final step by the reward model, specifically $\mathcal{R}(\mathbf{s}_{i},\mathbf{a}_{i})\triangleq R(\mathbf{x}_{T},c)$ if $i=T$, and 0 otherwise.
+In MixGRPO, we propose a sampling method that combines SDE and ODE.
+MixGRPO defines a time interval $S=[t_{l},t_{r})\subseteq[0,1)$, which corresponds to a subset of denoising timesteps, such that $0\leq l<r\leq T$ and $t_{i}=\frac{i}{T}$.
+We use SDE sampling within the interval $S$ and ODE sampling outside, while $S$ shifts along the denoising direction throughout the training process (See Figure 2 Left).
+MixGRPO restricts the stochastic exploration to the interval $S$, shortening the sequence length of the MDP to a subset $\Gamma_{\text{MixGRPO{}}}=(\mathbf{s}_{l},\mathbf{a}_{l},\mathbf{s}_{l+1},\mathbf{a}_{l+1},\ldots,\mathbf{s}_{r},\mathbf{a}_{r})$ and requires reinforcement learning (RL) optimization only within $S$:
+$\max_{\theta}\mathbb{E}_{\Gamma_{\text{MixGRPO}}\sim\pi_{\theta}}\bigg[$ $\sum_{t\in S}\bigg(\mathcal{R}\left(\mathbf{s}_{t},\mathbf{a}_{t}\right)$ (1)
+$-\beta D_{\text{KL}}\Big(\pi\left(\cdot|\mathbf{s}_{t}\right)||\pi_{\text{ref}}\left(\cdot|\mathbf{s}_{t}\right)\Big)\bigg)\bigg],$
+MixGRPO reduces training overhead while also enabling more concentrated gradient updates. Next, we derive the specific sampling form and optimization objective.
+For a deterministic reverse probability flow ODE [34], it takes the following form:
+$\frac{\mathrm{d}\mathbf{x}_{t}}{\mathrm{d}t}=f(\mathbf{x}_{t},t)-\frac{1}{2}g^{2}(t)\nabla_{\mathbf{x}_{t}}\log{q_{t}(\mathbf{x}_{t})},$ (2)
 where $q_{t}(\mathbf{x}_{t})$ represents the evolution process of the reverse probability distribution from 0 to $T$.
 $\nabla_{\mathbf{x}_{t}}\log{q_{t}(\mathbf{x}_{t})}$ is the score function at time $t$.
-According to the Fokker-Planck equation (Risken & Risken, 1996; Øksendal, 2003),  Song et al.
-(2020b) has demonstrated that Eq.
+According to the Fokker-Planck equation [28, 24],  [34] has demonstrated that Eq.
 (2) has the following equivalent probability flow SDE, which maintains the same marginal distribution at each time $t$:
-$\frac{\mathrm{d}\mathbf{x}_{t}}{\mathrm{d}t}=f(\mathbf{x}_{t},t)-g^{2}(t)\nabla_{\mathbf{x}_{t}}\log{q_{t}(\mathbf{x}_{t})}+g(t)\frac{\mathrm{d}\mathbf{w}}{\mathrm{d}t},\quad\mathbf{x}_{0}\sim q_{0}(\mathbf{x}_{0}).$ (3)
-In MixGRPO, we mix ODE and SDE for sampling, which has the same convergence as using only ODE sampling (a detailed proof in Appendix A). The specific form is as follows:
-[…]
-In particular, for Flow Matching (FM) (Lipman et al., 2022), especially the Rectified Flow (RF) (Liu et al., 2022), the sampling process can be viewed as a deterministic ODE:
+$\frac{\mathrm{d}\mathbf{x}_{t}}{\mathrm{d}t}=f(\mathbf{x}_{t},t)-g^{2}(t)\nabla_{\mathbf{x}_{t}}\log{q_{t}(\mathbf{x}_{t})}+g(t)\frac{\mathrm{d}\mathbf{w}}{\mathrm{d}t}.$ (3)
+In MixGRPO, we mix ODE and SDE for sampling, which has the same convergence as using only ODE sampling (a detailed proof in Supplementary). The specific form is as follows:
+$\mathrm{d}\mathbf{x}_{t}\!=\!\begin{cases}\left[f(\mathbf{x}_{t},t)\!-g^{2}(t)\mathbf{s}_{t}(\mathbf{x}_{t})\right]\mathrm{d}t+g(t)\mathrm{d}\mathbf{w},\!\!&\text{if}~t\in S,\\ \left[f(\mathbf{x}_{t},t)\!-\frac{1}{2}g^{2}(t)\mathbf{s}_{t}(\mathbf{x}_{t})\right]\mathrm{d}t,\!\!&\text{otherwise},\end{cases}$ (4)
+where we define the score function as $\mathbf{s}_{t}\ \triangleq\nabla_{\mathbf{x}_{t}}\log{q_{t}(\mathbf{x}_{t})}$.
+In particular, for Flow Matching (FM) [16], especially the Rectified Flow (RF) [19], the sampling process can be viewed as a deterministic ODE:
 $\frac{\mathrm{d}\mathbf{x}_{t}}{\mathrm{d}t}=\mathbf{v}_{t}.$ (5)
@@
 (5) is actually a special case of the Eq.
-(2) with $\mathbf{v}_{t}=f(\mathbf{x}_{t},t)-\frac{1}{2}g^{2}(t)\nabla_{\mathbf{x}_{t}}\log{q_{t}(\mathbf{x}_{t})}$.
-So we can derive the ODE-SDE hybrid sampling form for RF as follows:
+(2) with $\mathbf{v}_{t}=f(\mathbf{x}_{t},t)-\frac{1}{2}g^{2}(t)\mathbf{s}_{t}(\mathbf{x}_{t})$.
+So we can derive mixed ODE-SDE sampling form for RF as follows:
 […]
-In the RF framework, the model is used to predict the velocity field of the deterministic ODE, represented as $\mathbf{v}_{\theta}(\mathbf{x}_{t},t)=\frac{\mathrm{d}\mathbf{x}_{t}}{\mathrm{d}t}$.
-Following Liu et al.
-(2025a), the score function is represented as $\nabla_{\mathbf{x}_{t}}\log{q_{t}(\mathbf{x}_{t})}=-\frac{\mathbf{x}_{t}}{t}-\frac{1-t}{t}\mathbf{v}_{\theta}(\mathbf{x}_{t},t)$.
+In the RF framework, the model is used to predict the velocity field, represented as $\mathbf{v}_{\theta}(\mathbf{x}_{t},t)=\frac{\mathrm{d}\mathbf{x}_{t}}{\mathrm{d}t}$.
+Following [17], the score function is represented as $\mathbf{s}_{t}(\mathbf{x}_{t})=-\frac{\mathbf{x}_{t}}{t}-\frac{1-t}{t}\mathbf{v}_{\theta}(\mathbf{x}_{t},t)$.
 The $g(t)$ is represented as the standard deviation of the noise $g(t)=\sigma_{t}$.
@@
 Applying Euler-Maruyama discretization for SDE and Euler discretization for ODE, we build the final denoising process in MixGRPO:
+$\mathbf{x}_{t+\Delta t}=\begin{cases}\mathbf{x}_{t}+\boldsymbol{\mu}_{\theta}(\mathbf{x}_{t},t)\Delta t+\sigma_{t}\sqrt{\Delta t}\epsilon,&\text{if }t\in S,\\ \mathbf{x}_{t}+\mathbf{v}_{\theta}(\mathbf{x}_{t},t)\Delta t,&\text{otherwise},\end{cases}$ (7)
+where the SDE drift term for the sampling process is defined as $\boldsymbol{\mu}_{\theta}(\mathbf{x}_{t},t)\triangleq\mathbf{v}_{\theta}(\mathbf{x}_{t},t)+\frac{{\sigma^{2}_{t}}\left(\mathbf{x}_{t}+\left(1-t\right)\mathbf{v}_{\theta}(\mathbf{x}_{t},t)\right)}{2t}$.
+According to Eq.
+(1), we only need to optimize GRPO [32] at the time within the interval $S$ for $N$ samples in the group.
+For an unbiased advantage estimation in GRPO, we constrain the source of variation across images generated from a single prompt.
+Specifically, we fix the position of the SDE window, thereby sourcing all stochasticity from an identical timestep range in a GRPO group.
+This not only concentrates the gradients within this window but is also essential for preserving the unbiased nature of the optimization.
+The final training objective is represented as follows:
+$\mathcal{J}_{\text{MixGRPO}}(\theta)$ $=\mathbb{E}_{c\sim\mathcal{C},~\{\mathbf{x}^{i}_{T}\}^{N}_{i=0}\sim\pi_{\theta_{\text{old}}}(\cdot|c)}\left[\frac{1}{N}\sum^{N}_{i=1}\frac{1}{|S|}\sum_{t\in S}\right$.
+(8)
+$\qquad\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\left.\text{min}\bigg(r^{i}_{t}(\theta)A^{i},\text{clip}\Big(r^{i}_{t}(\theta),1\!-\!\varepsilon,1\!+\!\varepsilon\Big)A^{i}\bigg)-\beta\mathcal{J}_{\text{KL}}\right],$
+where $\varepsilon$ is the clipping hyperparameter, $r^{i}_{t}(\theta)$ is the policy ratio, $A^{i}$ is the advantage score and $\mathcal{J}_{KL}$ is the KL-regularization, defined as follows:
+$r^{i}_{t}(\theta)$ $=\frac{q_{\theta}(\mathbf{x}_{t+\Delta t}|\mathbf{x}_{t},c)}{q_{\theta_{\text{old}}}({\mathbf{x}_{t+\Delta t}|\mathbf{x}_{t},c})},$ (9)
+$A^{i}$ $=\frac{R\left(\mathbf{x}^{i}_{T},c\right)-\text{mean}\left(\{R\left(\mathbf{x}^{i}_{T},c\right)\}^{N}_{i=1}\right)}{\text{std}\left(\{R\left(\mathbf{x}^{i}_{T},c\right)\}^{N}_{i=1}\right)},$
 […]
-According to Eq. (1), we only need to optimize GRPO (Shao et al., 2024) at the time within the interval $S$ for $N$ samples in the group. The final training objective is represented as follows:
-$\mathcal{J}_{\text{MixGRPO}}(\theta)$ $=\mathbb{E}_{c\sim\mathcal{C},~\{\mathbf{x}^{i}_{T}\}^{N}_{i=0}\sim\pi_{\theta_{\text{old}}}(\cdot|c)}$ (8)
-$\left[\frac{1}{N}\sum^{N}_{i=1}\frac{1}{|S|}\sum_{t\in S}\text{min}\left(r^{i}_{t}(\theta)A^{i},\text{clip}\left(r^{i}_{t}(\theta),1-\varepsilon,1+\varepsilon\right)A^{i}\right)\right],$
-where $r^{i}_{t}(\theta)$ is referred to as the policy ratio and $A^{i}$ is the advantage score. We set $\varepsilon=0.0001$,
-$r^{i}_{t}(\theta)=\frac{q_{\theta}(\mathbf{x}_{t+\Delta t}|\mathbf{x}_{t},c)}{q_{\theta_{\text{old}}}({\mathbf{x}_{t+\Delta t}|\mathbf{x}_{t},c})}\quad\text{and}\quad A^{i}=\frac{R\left(\mathbf{x}^{i}_{T},c\right)-\text{mean}\left(\{R\left(\mathbf{x}^{i}_{T},c\right)\}^{N}_{i=1}\right)}{\text{std}\left(\{R\left(\mathbf{x}^{i}_{T},c\right)\}^{N}_{i=1}\right)},$ (9)
-It is important to note that, we have dropped the KL Loss.
-Although KL Loss can mitigate reward hacking to some extent (Liu et al., 2025a), inspired by yifan123 (2025), we use hybrid inference at test time, which can significantly address the reward hacking issue (See Appendix B).
-MixGRPO reduces NFE of $\pi_{\theta}$ compared to all-timesteps optimization.
+MixGRPO reduces the NFE of $\pi_{\theta}$ and optimized timesteps count compared to global SDE sampling.
 However, the NFE of $\pi_{\theta_{\text{old}}}$ is not reduced, as complete inference is required to obtain the final image for reward calculation.
 In Section 3.3, we will introduce the use of higher-order ODE solvers, which also reduce the NFE of $\pi_{\theta_{\text{old}}}$ leading to further speedup.
-In summary, the mixed ODE-SDE sampling significantly reduces overhead while ensuring the introduction of randomness, allowing for optimization by GRPO.
+In summary, the mixed ODE-SDE sampling significantly streamlines the MDP, enhancing training efficiency by lowering optimization overhead and enabling more focused gradient updates.
+Algorithm 1 MixGRPO Training Process
+1:initial policy model $\pi_{\theta}$; reward models $\{R_{k}\}_{k=1}^{K}$; prompt dataset $\mathcal{C}$; total sampling steps $T$; number of samples per prompt NN;
+2:sliding window $W(l)$, window size $w$, shift interval $\tau$, window stride $s$
+3:Init left boundary of $W(l)$: $l\leftarrow 0$
+4:for training iteration $m=1$ to $M$ do
+5:  Sample batch prompts $\mathcal{C}_{b}\sim\mathcal{C}$
+6:  Update old policy model: $\pi_{\theta_{\text{old}}}\leftarrow\pi_{\theta}$
+7:  for each prompt $\mathbf{c}\in\mathcal{C}_{b}$ do
+8:    Init the same noise $\mathbf{x_{0}}\sim\mathcal{N}(0,\mathbf{I})$
+9:    for generate $i$-th image from $i=1$ to N do
+10:     for sampling timestep $t=0$ to T1T-1 do
+11:      if $t\in W(l)$ then
+12:         Use SDE Sampling to get $\mathbf{x}^{i}_{t+1}$
+13:      else
+14:         Use ODE Sampling to get $\mathbf{x}^{i}_{t+1}$
+15:      end if
+16:     end for
+17:    end for
+18:    for $i$-th image from $i=1$ to N do
+19:     Calculate advantages: $A_{i}\leftarrow\sum_{k=1}^{K}\frac{R(\mathbf{x}^{i}_{T},\mathbf{c})^{i}_{k}-\mu_{k}}{\sigma_{k}}$
+20:    end for
+21:    for optimization timestep $t\in W(l)$ do
+22:     Update policy model: $\theta\leftarrow\theta+\eta\nabla_{\theta}\mathcal{J}$
+23:    end for
+24:  end for
+25:  if $m\bmod\tau~\text{is}~0$ then $\triangleright$ move sliding window
+26:    $l\leftarrow\min(l+s,~T-w)$
+27:  end if
+28:end for
 3.2 Sliding Window as Optimization Scheduler
 Figure 3: Qualitative comparison. MixGRPO achieve superior performance in semantics, aesthetics and text-image alignment.
-In fact, the interval $S$ can be non-fixed during the training process.
 In this section, we will introduce the sliding window to describe the movement of $S$, which leads to a significant improvement in the quality of the generated images.
-Along the denosing time-steps $\{0,1,\ldots,T-1\}$, MixGRPO defines a SDE sliding window $W(l)$ and optimization is only employed at the steps within $W(l)$.
+Along the denosing time-steps $\{0,1,\ldots,T-1\}$, MixGRPO defines a SDE sliding window $W(l)$ and optimization is only employed at the timesteps within $W(l)$.
 $W(l)=\{t_{l},t_{l+1},\dots,t_{l+w-1}\},\quad l\leq T-w,$ (10)
@@
 The left boundary $l$ of the sliding window moves as the training progresses.
-In the experiments, we found that the window size $w$, shift interval $\tau$, and window stride $s$ are all crucial hyperparameters.
-Through ablation studies (See Experiment 4.4.1), we identified the optimal settings.
-When the total sampling steps $T=25$, the best performance is achieved with $w=4$, $\tau=25$ and $s=1$.
+The movement strategy of the sliding window is governed by the following parameters.
+$\tau$ is the shift interval, which specifies the number of training steps between each window shift.
+$s$ is the window stride, which indicates the number of denoising timesteps that the left boundary $l$ advances during each shift.
 The detailed sliding window strategy and MixGRPO algorithm can be found in Algorithm 1.
-Restricting the use of SDE sampling within the sliding window not only ensures the diversity of the generated images but also allows the model to concentrate on optimizing the flow within that window.
-Moving along the denoising direction represents the stochasticity of the probability flow from strong to weak, as illustrated in Figure 2.
-This is essentially a greedy strategy and is similar to the RL that assigns discount factors to process rewards (Pitis, 2019; Amit et al., 2020; Hu et al., 2022b), which gives greater importance to rewards derived from a larger search space in the earlier process.
+In Section 4.4.1, we conduct thorough ablation studies to validate these parameters.
+In our MixGRPO, the denoising process inherently follows a stochastic-to-deterministic trajectory: moving along the probability flow corresponds to a transition from high to low stochasticity, as illustrated in Figure 2 (right).
+Our sliding-window approach operationalizes this principle by creating an implicit training curriculum.
+This is analogous to reward shaping in RL [26, 2, 10], where discount factors prioritize rewards from the most critical steps in a process.
+For diffusion and flow-matching models, this curriculum directs the GRPO algorithm to first optimize low-SNR regions—where dynamics are highly stochastic—before shifting focus to the more deterministic, high-SNR regions for fine-grained refinement.
+This coarse-to-fine strategy stabilizes the initial, more volatile stages of training and enables the model to learn a robust global structure before perfecting finer details.
 The experimental results of different movement strategies in Table 4 demonstrate the validity of this intuition.
-Not only does the progressive strategy outperform random selection, but we also found that even when the sliding window is not moved (frozen), meaning only the earlier timesteps are optimized, MixGRPO can still yield good results, particularly in terms of ImageReward (Xu et al., 2023) and UnifiedReward (Wang et al., 2025).
+Not only does the progressive strategy outperform random selection, but our findings also indicate that MixGRPO can achieve strong performance even with a static (frozen) window.
+In this setting, where optimization is exclusively applied to the initial timesteps, our method still surpasses DanceGRPO.
+This is because the initial timesteps in the generative process, which restores an image from Gaussian noise, account for a more substantial degree of denoising than other timesteps [11, 4].
+Initial timesteps introduces greater stochasticity into the final output, thereby increasing the likelihood of generating images that align with the preferences of a reward model and ultimately yielding a higher optimization payoff.
 Based on this intuition, we also proposed an exponential decay strategy as follows, allowing $\tau$ to decrease as the window moves, enabling the model to avoid excessive optimization in smaller search spaces.
@@
 The exponential function $\exp(x)$ calculates $e^{x}$, while the Rectified Linear Unit $\text{ReLU}(x)$ is defined as $\max(0,x)$.
-Table 5 shows that the exponential decay strategy can achieve better results in terms of Pick Score (Kirstain et al., 2023) and ImageReward (Xu et al., 2023).
+Table 4 shows that the exponential decay strategy can achieve better results in terms of Pick Score [12] and ImageReward [41].
 This may be because the model focuses on the earlier stages of denoising, which can lead to more significant high-level changes, precisely what the human preference alignment reward model emphasizes.
 3.3 Trade-off Between Overhead and Performance
-MixGRPO employs SDE sampling within the sliding window and ODE sampling outside of it, allowing the use of higher-order ODE solvers to accelerate GRPO training-time sampling.
-The timesteps that utilize ODE sampling are divided into those before and after the sliding window.
-The timesteps after the sliding window solely influence the reward calculation, whereas the timesteps before the window affect both the reward and contribute to cumulative errors in the policy ratio computing.
-Therefore, we focus exclusively on the acceleration of the timesteps after the window.
-Gao et al.
-(2024) has demonstrated the equivalence between the ODE sampling of flow matching models (FM) and DDIM, and Section 3.1 has also shown that diffusion probabilistic models (DPM) and FM share the same ODE form during the denoising process.
-Therefore, the higher-order ODE solvers e.g., DPM-Solver Series (Lu et al., 2022a; b; Zheng et al., 2023), UniPC (Zhao et al., 2023) designed for DPM sampling acceleration are also applicable to FM.
-We have reformulated DPM-Solver++ (Lu et al., 2022b) to apply it in the FM framework for ODE sampling acceleration and released detailed derivations in Appendix C.
+Unlike DanceGRPO [42], which relies on global SDE sampling, MixGRPO employs a mixed ODE-SDE method, allowing the use of higher-order ODE solvers to accelerate GRPO training-time sampling.
+However, accelerating the ODE phase before the SDE window amplifies the solver’s numerical errors due to the window’s stochasticity, severely degrading image quality and corrupting the reward signal.
+In contrast, accelerating the ODE phase after the window provides a significant speed-up without compromising the image fidelity required for reliable reward evaluation.
+Therefore, MixGRPO exclusively accelerates the post-window ODE timesteps.
+[7] has demonstrated the equivalence between the ODE sampling of flow matching models (FM) and DDIM, and Section 3.1 has also shown that diffusion probabilistic models (DPM) and FM share the same ODE form during the denoising process.
+Therefore, the higher-order ODE solvers e.g., DPM-Solver Series [21, 22, 47], UniPC [46] designed for DPM sampling acceleration are also applicable to FM.
+We have reformulated DPM-Solver++ [22] to apply it in the FM framework for ODE sampling acceleration and released detailed derivations in Supplementary.
 By applying higher-order solvers, we achieve acceleration in the sampling of $\pi_{\theta_{\text{old}}}$ during GRPO training, which is essentially a balance between overhead and performance.
@@
 Ultimately, we introduced DPM-Solver++ and adopted both progressive and frozen sliding window strategies, proposing MixGRPO-Flash and MixGRPO-Flash*.
-A detailed description of the algorithm can be found in Appendix E.
-These approaches achieve a greater degree of acceleration compared to MixGRPO, while also outperforming DanceGRPO in terms of human preference alignment performance.
+A detailed description of the algorithm can be found in Supplementary.
+These faster versions offer more substantial acceleration relative to base version of our MixGRPO, while simultaneously surpassing DanceGRPO in their alignment with human preferences.
 Table 1: Comparison results for overhead and performance.
@@
 4.1 Experiment Setup
-Dataset  We conduct experiments using the prompts provided by the HPDv2111https://huggingface.co/datasets/ymhao/HPDv2 dataset, which is the official dataset for the HPS-v2 benchmark (Wu et al., 2023).
+Dataset  We conduct experiments using the prompts provided by the HPDv2 dataset, which is the official dataset for the HPS-v2 benchmark [40].
 The training set contains 103,700 prompts; in fact, MixGRPO achieved good human preference alignment results after training one epoch with only 9,600 prompts.
@@
 The prompts are diverse, encompassing four styles: “Animation”, “Concept Art”, “Painting”, and “Photo”.
-Model  Following DanceGRPO (Xue et al., 2025), we use FLUX.1 Dev (Labs, 2024) as the base model, which is an advanced text-to-image model based on flow matching.
-Overhead Evaluation  For the evaluation of overhead, we use two metrics: the number of function evaluations (NFE) (Lu et al., 2022a) and the time consumption per iteration during training.
+Model  Following DanceGRPO [42], we use FLUX.1 Dev [13] as the base model, which is an advanced text-to-image model based on flow matching.
+Overhead Evaluation  For the evaluation of overhead, we use two metrics: the number of function evaluations (NFE) [21] and the time consumption per iteration during training.
 The NFE is divided into $\text{NFE}_{\pi_{\theta_{\text{old}}}}$ and $\text{NFE}_{\pi_{\theta}}$.
@@
 Additionally, the average training time per GRPO iteration provides a more accurate reflection of the acceleration effect.
-Performance Evaluation  We used four multiple reward models in conjunction for GRPO, namely HPS-v2.1 (Wu et al., 2023), Pick Score (Kirstain et al., 2023), ImageReward (Xu et al., 2023) and Unified Reward Wang et al.
-(2025), both as reward guidance during training and as evaluation metrics.
+Performance Evaluation  We used four multiple reward models in conjunction for GRPO, namely HPS-v2.1 [40], Pick Score [12], ImageReward [41] and Unified Reward [37], both as reward guidance during training and as evaluation metrics.
 These metrics are all based on human preferences but emphasize different aspects.
-For example, ImageReward (Xu et al., 2023) highlights image-text alignment and fidelity, while Unified Reward (Wang et al., 2025) concentrates on semantics.
-DanceGRPO Xue et al.
-(2025) also demonstrates that using multiple reward models can achieve better results.
-To validate the robustness of MixGRPO, we also followed DanceGRPO and conducted additional comparisons using HPS-v2.1 as a single reward, and combining HPS-v2.1 (Wu et al., 2023) and CLIP Score (Radford et al., 2021) as multi-rewards.
+For example, ImageReward [41] highlights image-text alignment and fidelity, while Unified Reward [37] concentrates on semantics.
+DanceGRPO [42] also demonstrates that using multiple reward models can achieve better results.
+To validate the robustness of MixGRPO, we also followed DanceGRPO and conducted additional comparisons using HPS-v2.1 as a single reward, and combining HPS-v2.1 [40] and CLIP Score [27] as multi-rewards.
 4.2 Implementation Details
@@
 Furthermore, when multiple reward models are jointly trained with, each reward is assigned equal weight.
-For training, all experiments are conducted on 32 Nvidia GPUs with a batch size of 1 and a maximum of 300 iterations.
-We use AdamW (Loshchilov & Hutter, 2017) as the optimizer with a learning rate of 1e-5 and a weight decay coefficient of 0.0001.
-Mixed precision training is used with the bfloat16 (bf16) format, while the master weights are maintained at full precision (fp32).
-4.3 Main Experiments
-In the main experiment, the four human-preference-based rewards were aggregated according to advantages, as shown in Algorithm 1.
-We evaluated the overhead and performance of MixGRPO in comparison with DanceGRPO, with the results presented in Table 1.
-The official DanceGRPO uses $\text{NFE}_{\pi_{\theta}}=14$; however, for fairness, we also tested DanceGRPO with $\text{NFE}_{\pi_{\theta}}=4$.
+For training, all experiments are conducted on 32 Nvidia GPUs with a batch size per gpu of 1 and a maximum of 300 iterations.
+We use AdamW [20] as the optimizer with a learning rate of 1e-5 and a weight decay coefficient of 0.0001.
+Mixed precision training is used with the bfloat16 (bf16) format, while the master weights are maintained at fp32 precision.
+4.3 Comparison with existing methods
+To validate the generality, effectiveness, and efficiency of our proposed method, we conducted comprehensive experiments across two distinct frameworks: Dance-GRPO and Flow-GRPO.
+In our experiments, we employed multiple reward models for guidance, as described in Algorithm 1.
+We first evaluated the overhead and performance of MixGRPO in comparison with DanceGRPO, with the results presented in Table 1.
+The official DanceGRPO uses $\text{NFE}_{\pi_{\theta}}=14$; For fairness, we also tested DanceGRPO with $\text{NFE}_{\pi_{\theta}}=4$.
 For MixGRPO-Flash, we evaluated both the progressive and frozen strategies, and to ensure fairness, we also applied the frozen strategy to DanceGRPO.
-We selected multiple scene prompts to visualize the results for FLUX.1 Dev, the officially configured DanceGRPO, and MixGRPO, as shown in Figure 3.
-It can be observed that MixGRPO achieved the best results in terms of semantics, aesthetics, and text-image alignment.
-Figure 5 shows the comparison results of DanceGRPO, MixGRPO, and MixGRPO-Flash with $\text{NFE}_{\pi_{\theta}}=4$.
-It can be observed that under the same overhead conditions, MixGRPO achieved better results compared to DanceGRPO.
-Additionally, MixGRPO-Flash enables accelerated sampling of $\pi_{\theta_{\text{old}}}$, and as the overhead decreases, the quality of the generated images still maintains a strong alignment with human preferences.
-Following DanceGRPO (Xue et al., 2025), we also trained and evaluated the model using a single reward model, e.g., HPSv2.1 and two reward models, e.g., HPSv2.1 and CLIP Score on the HPDv2 dataset (Wu et al., 2023).
+The visual results are presented in Figure 3.
+We observe that MixGRPO achieves the best performance in terms of semantics, aesthetics, and text-image alignment.
+For a direct comparison at $\text{NFE}_{\pi_{\theta}}=4$, the qualitative results are shown in Figure 4.
+confirms the robustness of acceleration in MixGRPO-Flash, demonstrating the model’s ability to preserve the aesthetic quality and semantic alignment while speeding up the process.
+Following DanceGRPO [42], we also trained and evaluated the model using a single reward model, e.g., HPSv2.1 and two reward models, e.g., HPSv2.1 and CLIP Score on the HPDv2 dataset [40].
 The results (See Table 2) demonstrate that MixGRPO achieves the best performance on both in-domain and out-of-domain rewards, whether using a single or multiple reward models.
-The visualized results are displayed in Figure 6 and Figure 7 of Appendix F.
-To demonstrate the robustness of the MixGRPO, we followed Flow-GRPO (Liu et al., 2025a) and applied the Low-Rank Adaptation (LoRA) (Hu et al., 2022a) method on the Stable Diffusion 3.5 (SD3.5) (Esser et al., 2024).
-We used HPS-v2.1 (Wu et al., 2023), Pick Score (Kirstain et al., 2023), and ImageReward (Xu et al., 2023) as multi-rewards, comparing Flow-DPO (offline/online) (Liu et al., 2025b), and Flow-GRPO (Liu et al., 2025a).
-The results in Table 3 showed that MixGRPO is well-suited for the LoRA, significantly reducing training overhead while outperforming both Flow-DPO and Flow-GRPO in terms of speed and performance.
-The visualization results can be found in Figure 8 of Appendix F.
-Table 2: Comparison results demonstrate that MixGRPO achieves the best performance on both in-domain and out-of-domain rewards.
+The visualized results are displayed in Supplementary.
+Table 2: Comparison with DanceGRPO. The results demonstrate that MixGRPO achieves the best performance on both in-domain and out-of-domain rewards.
 Reward Model Method In Domain Out-of-Domain
@@
 MixGRPO 0.349 0.415 0.229 1.416 3.430
-Table 3: Comparison results demonstrate that MixGRPO outperforms Flow-DPO and Flow-GRPO in training efficiency and performance.
+Table 3: Comparison with FlowDPO and FLow-GRPO. The results demonstrate that MixGRPO outperforms others in training efficiency and performance.
 Model RL Method $\text{NFE}_{\pi_{\theta_{\text{old}}}}$ $\text{NFE}_{\pi_{\theta}}$ HPSv2.1 Pick Score ImageReward
@@
 MixGRPO 10 4 0.3416 0.2360 1.4854
+Methodologically, Flow-GRPO [17] and DanceGRPO [42] are conceptually similar, differing only in minor aspects such as their sampling formulations and the use of KL-regularization.
+Therefore, to ensure a fair and direct comparison, we benchmarked our method against their respective officially released models.
+To demonstrate the robustness of the MixGRPO across distinct foundational models, we also followed Flow-GRPO [17] and applied the Low-Rank Adaptation (LoRA) [9] on the Stable Diffusion 3.5(SD3.5) [4].
+We used HPS-v2.1 [40], Pick Score [12], and ImageReward [41] as multi-rewards, comparing Flow-DPO (offline/online) [18], and Flow-GRPO [17].
+The results in Table 3 showed that MixGRPO is well-suited for the LoRA, significantly reducing training overhead while outperforming both Flow-DPO and Flow-GRPO in terms of speed and performance.
+The visualization results can be found in Supplementary.
 4.4 Ablation Experiments
 4.4.1 Sliding Window Hyperparamters
-As introduced in Section 3.2, the moving strategy, shift interval $\tau$, window size $w$ and window stride $s$ are all important parameters of the sliding window.
+Table 4: Comparison for moving strategies.
+Strategy Interval Schedule HPS-v2.1 Pick Score ImageReward Unified Reward
+Frozen / 0.354 0.234 1.580 3.403
+Random Constant 0.365 0.237 1.513 3.388
+Progressive Decay (Linear) 0.365 0.235 1.566 3.382
+Decay (Exp) 0.360 0.239 1.632 3.416
+Constant 0.367 0.237 1.629 3.418
+Table 5: Comparison for different shift interval $\tau$.
+$\tau$ HPS-v2.1 Pick Score ImageReward Unified Reward
+15 0.366 0.237 1.509 3.403
+20 0.366 0.238 1.610 3.411
+25 0.367 0.237 1.629 3.418
+30 0.350 0.229 1.589 3.385
+As introduced in Section 3.2, the moving strategy, shift interval $\tau$, window size $w$ and window stride $s$ are parameters introduced by the SDE sliding window.
 We conducted ablation experiments on each of them.
@@
 For the window stride $s$, we found through experimentation that $s=1$ is the optimal choice, as shown in Table 7.
-Table 4: Comparison for moving strategies.
-Strategy
-Interval
-Schedule
-HPS-v2.1 Pick Score ImageReward Unified Reward
-Frozen / 0.354 0.234 1.580 3.403
-Random Constant 0.365 0.237 1.513 3.388
-Progressive Decay (Linear) 0.365 0.235 1.566 3.382
-Decay (Exp) 0.360 0.239 1.632 3.416
-Constant 0.367 0.237 1.629 3.418
-Table 5: Comparison for shift interval $\tau$.
-$\tau$ HPS-v2.1 Pick Score ImageReward Unified Reward
-15 0.366 0.237 1.509 3.403
-20 0.366 0.238 1.610 3.411
-25 0.367 0.237 1.629 3.418
-30 0.350 0.229 1.589 3.385
-Table 6: Comparison for window size $w$
+In practice, the optimal settings are applicable not only to Flux [13] but also to Stable Diffusion 3.5 [4].
+Moreover, the model can maintain its optimal generation performance within a controllable range of parameter adjustments.
+This demonstrates the robustness of the sliding-window optimization scheduling strategy.
+Table 6: Comparison for different window size $w$.
 $w$ $\text{NFE}_{\pi_{\theta}}$ HPS-v2.1 Pick Score ImageReward Unified Reward
@@
 6 6 0.370 0.238 1.547 3.398
-Table 7: Comparison for window stride $s$
+Table 7: Comparison for different window stride $s$.
 $s$ HPS-v2.1 Pick Score ImageReward Unified Reward
@@
 4 0.368 0.238 1.575 3.407
-4.4.2 High Order ODE Solver
-MixGRPO enables the possibility of accelerating ODE sampling with high-order ODE solvers by combining SDE and ODE sampling methods.
-We first conducted ablation experiments on the order of the solver, using DPM-Solver++ (Lu et al., 2022b) as the high-order solver with the progressive strategy.
-The results, as shown in Table 9, indicate that the second-order mid-point method is the optimal setting.
-Then, as described in Section 3.3, we compared two acceleration approaches.
-One is MixGRPO-Flash, which utilizes the progressive window moving strategy.
-The other is MixGRPO-Flash*, which employs the frozen moving strategy.
-They all achieve a balance between overhead and performance by reducing the number of ODE sampling steps after the sliding window.
-However, in practice, MixGRPO-Flash requires the window to move throughout the training process, which results in a shorter ODE portion being accelerated.
-Consequently, the acceleration effect of MixGRPO-Flash, on average, is not as pronounced as that of MixGRPO-Flash*.
-Table 8: Comparison of performance across different-order solvers. The second-order Midpoint method achieves the best performance.
+Table 8: Comparison of performance across different solvers. The second-order Midpoint method achieves the best performance.
 Order Solver Type HPS-v2.1 Pick Score ImageReward Unified Reward
@@
 3 / 0.359 0.234 1.512 3.387
-Table 9: Comparison of different sampling steps for $\pi_{\theta_{\text{old}}}$. MixGRPO-Flash* achieves good performance even with few steps.
+4.4.2 High Order ODE Solver
+MixGRPO-Flash, which incorporates a high-order ODE solver to accelerate the sampling process of the reference model, achieves an effective trade-off between speed and performance.
+For MixGRPO-Flash, we first conducted ablation experiments on the order of the solver, using DPM-Solver++ [22] as the high-order solver with the progressive strategy.
+The results, as shown in Table 8, indicate that the second-order mid-point method is the optimal setting, optimizing the most human preference alignment metrics while simultaneously accelerating the process.
+Then we compared two acceleration approaches.
+One is MixGRPO-Flash, and the other is MixGRPO-Flash*.
+Both utilize a second-order ODE solver for acceleration, but they differ in their sliding-window moving strategies.
+The quantitative results are presented in Table 9.
+MixGRPO-Flash requires the window to move throughout the training process, resulting in a smaller portion of the ODE being accelerated compared to MixGRPO-Flash*.
+Consequently, MixGRPO-Flash* not only achieves a higher degree of acceleration for the reference model but also yields superior results in the ImageReward [41] and Unified Reward [37] metrics.
+Figure 4: Qualitative results of MixGRPO-Flash. As the training overhead is reduced, generated images maintain high quality.
+Table 9: Comparison about different sampling steps of the reference model. MixGRPO-Flash and MixGRPO-Flash* all achieve a balance between overhead and performance.
 Method Sampling Overhead Human Preference Alignment
@@
 8 3.789 0.357 0.232 1.624 3.402
+5 Limitation
+A key limitation of GRPO is that its performance ceiling is inherently tied to the capability of the reward model.
+Specifically, when a reward model is not sufficiently powerful to provide a comprehensive assessment of image quality, the GRPO algorithm is prone to reward hacking, particularly in the later stages of training.
+As confirmed by several studies [39, 38, 23], reward hacking is fundamentally an issue stemming from the limitations of the reward model and is less attributable to the RL algorithm itself.
+In this context, the primary focus of MixGRPO is not to overcome the reward model’s limitations, but rather to explore how to achieve faster convergence and superior performance under the guidance of existing, imperfect reward models.
+Looking ahead, while the ultimate performance is contingent upon the reward model’s quality, we plan to explore and train more powerful reward models as a direction for future work.
 6 Conclusion
-Although GRPO (Shao et al., 2024) has seen significant success in the language modality, it is still in the early stages of progress in vision (Tong et al., 2025; Xue et al., 2025; Liu et al., 2025a).
-Existing flow-based GRPO faces challenges such as low sampling efficiency and slow training.
-To address these issues, we proposed MixGRPO, a novel training framework that combines SDE and ODE sampling.
-This hybrid approach allows for focused optimization on the SDE sampling flow component, reducing complexity while ensuring accurate reward computation.
-Inspired by the decay factor in reinforcement learning (Hu et al., 2022b), we introduce a sliding window strategy for scheduling the optimized denoising steps.
-Experimental results confirm the effectiveness of our approach in both single-reward and multi-reward settings.
-Additionally, MixGRPO decouples the denoising stages for optimization and reward computation, enabling acceleration of the latter with high-order solvers.
-We further propose MixGRPO-Flash, which balances overhead and performance.
-We hope MixGRPO will inspire deeper research into post-training for image generation, contributing to the advancement of Artificial General Intelligence (AGI).
-6 Acknowledgement
-The authors thank Yiming Cheng (M.Sc., Tsinghua University, 13306339695@163.com) for his valuable feedback on the theoretical derivations and mathematical expressions.
-Appendix A Proof of convergence for Mixed ODE-SDE sampling
+In this work, we have presented MixGRPO, a novel hybrid ODE-SDE framework for improving GRPO training efficiency and performance.
+We have proposed a strategy to confine optimization to a dynamic stochastic interval managed by a sliding window.
+This guides the optimization process from broad exploration to fine-grained refinement, thus enhancing performance.
+Experiments demonstrate that MixGRPO has achieved superior performance in both single-reward and multi-reward settings while substantially reducing overhead.
+Furthermore, we have presented MixGRPO-Flash, a variant offering a flexible trade-off between performance and computational cost.
+Supplementary Material
+7 Proof of Convergence for Mixed ODE-SDE Sampling
 To prove that the mixed ODE-SDE sampling method in Eq.
 (7) has the same convergence as Eq.
-(2), which uses only ODE sampling, referencing Song et al.
-(2020b), we approach this from the perspective of distribution evolution, where the distribution at each time step, e.g., $\frac{\partial{q_{t}(\mathbf{x})}}{\partial{t}}$ must be the same.
+(2), which uses only ODE sampling, referencing [34], we approach this from the perspective of distribution evolution, where the distribution at each time step, e.g., $\frac{\partial{q_{t}(\mathbf{x})}}{\partial{t}}$ must be the same.
 Let the interval for SDE be defined as $S=\left[t_{l},t_{r}\right)\in\left[0,1\right)$.
@@
 $\mathrm{d}\mathbf{x}=[f(\mathbf{x},t)-g^{2}(t)\nabla_{\mathbf{x}}\log{q_{t}(\mathbf{x})}]\mathrm{d}t+g(t)\mathrm{d}\mathbf{w},\quad t\in S.$ (12)
-The marginal probability density $q_{t}(\mathbf{x}_{t})$ evolves according to Kolmogorov’s equation (Fokker-Planck equation) (Øksendal, 2003)
-$\frac{\partial q_{t}(\mathbf{x})}{\partial t}=-\nabla_{\mathbf{x}}\cdot\big[\big(f(\mathbf{x},t)-g^{2}(t)\nabla_{\mathbf{x}}\log q_{t}(\mathbf{x})\big)q_{t}(\mathbf{x})\big]+\frac{1}{2}g^{2}(t)\nabla^{2}_{\mathbf{x}}q_{t}(\mathbf{x})$ (13)
+The marginal probability density $q_{t}(\mathbf{x}_{t})$ evolves according to Kolmogorov’s equation (Fokker-Planck equation) [24]
+$\frac{\partial q_{t}(\mathbf{x})}{\partial t}=$ $-\nabla_{\mathbf{x}}\cdot\big[\big(f(\mathbf{x},t)-g^{2}(t)\nabla_{\mathbf{x}}\log q_{t}(\mathbf{x})\big)q_{t}(\mathbf{x})\big]$ (13)
+$+\frac{1}{2}g^{2}(t)\nabla^{2}_{\mathbf{x}}q_{t}(\mathbf{x})$
 […]
-$\frac{\partial q_{t}(\mathbf{x})}{\partial t}$ $=-\nabla_{\mathbf{x}}\cdot\big[f(\mathbf{x},t)q_{t}(\mathbf{x})-g^{2}(t)\nabla_{\mathbf{x}}q_{t}(\mathbf{x})\big]+\frac{1}{2}g^{2}(t)\nabla^{2}_{\mathbf{x}}q_{t}(\mathbf{x})$
+$\frac{\partial q_{t}(\mathbf{x})}{\partial t}$ $=-\nabla_{\mathbf{x}}\cdot\big[f(\mathbf{x},t)q_{t}(\mathbf{x})-g^{2}(t)\nabla_{\mathbf{x}}q_{t}(\mathbf{x})\big]$ (14)
+$\qquad~+\frac{1}{2}g^{2}(t)\nabla^{2}_{\mathbf{x}}q_{t}(\mathbf{x})$
 $=-\nabla_{\mathbf{x}}\cdot[f(\mathbf{x},t)q_{t}(\mathbf{x})-\frac{1}{2}g^{2}(t)\nabla_{\mathbf{x}}q_{t}(\mathbf{x})]$
-$=-\nabla_{\mathbf{x}}\cdot\big[\underbrace{\big(f(\mathbf{x},t)-\frac{1}{2}g^{2}(t)\nabla_{\mathbf{x}}\log q_{t}(\mathbf{x})\big)}_{f_{\text{ODE}}(\mathbf{x},t)}q_{t}(\mathbf{x})\big].$ (14)
-The Eq. (A) is indeed the Fokker-Planck equation of the ODE Eq. (2). Therefore, within the interval $S$, the distribution evolution of SDE and ODE sampling is consistent.
-Appendix B Hybrid Inference for Solving Reward Hacking
-To address the reward hacking issue, we employ hybrid inference to sample with both the raw and post-trained models, and introduce the hybrid percent $p_{\text{mix}}$.
-This means that the initial $p_{\text{mix}}T$ denoising steps are sampled by the model trained with GRPO, while the remaining denoising process is finished by the original FLUX model (Black et al., 2023).
-Table 10 and Figure 4 respectively illustrate the changes in performance and images as $p_{\text{mix}}$ increases under the multi-rewards training scenario.
-The experimental results demonstrate that $p_{\text{mix}}=80\%$ is an optimal empirical value that effectively mitigates hacking while maximizing alignment with human preferences.
-Table 10: Comparison with different hybrid inference percentages
-$p_{\text{mix}}$ HPS-v2.1 Pick Score ImageReward Unified Reward
-0% 0.313 0.226 1.089 3.369
-20% 0.342 0.233 1.372 3.386
-40% 0.356 0.235 1.539 3.395
-60% 0.362 0.236 1.598 3.407
-80% 0.366 0.238 1.610 3.411
-100% 0.369 0.238 1.607 3.378
-Figure 4: Qualitative comparison with different hybrid inference percentages
-Appendix C DPM-Solver++ for Recitified Flow
+$=-\nabla_{\mathbf{x}}\!\cdot\!\big[\!\underbrace{\big(f(\mathbf{x},t)\!-\!\frac{1}{2}g^{2}(t)\nabla_{\mathbf{x}}\log q_{t}(\mathbf{x})\big)}_{f_{\text{ODE}}(\mathbf{x},t)}q_{t}(\mathbf{x})\!\big].$
+The Eq. (14) is indeed the Fokker-Planck equation of the ODE Eq. (2). Therefore, within the interval $S$, the distribution evolution of SDE and ODE sampling is consistent.
+8 DPM-Solver++ for Recitified Flow
 For clarity and to avoid ambiguity between continuous time and discrete steps, we adopt the following notation in this section.
@@
 The continuous time corresponding to step $i$ is denoted by $t_{i}=\frac{i}{T}\in[0,1)$.
-The DPM-Solver++ algorithm (Lu et al., 2022b) is originally designed for the $\mathbf{x}_{0}\text{-}prediction$ diffusion model (Rombach et al., 2022), where the model outputs the denoised feature $\mathbf{x}_{0}$ based on the noisy feature $\mathbf{x}_{t_{i}}$, the time condition $t_{i}$ and the text condition $c$.
-According to the definition of Rectified Flow (RF) (Liu et al., 2022), there is the following transfer equation:
+The DPM-Solver++ algorithm [22] is originally designed for the $\mathbf{x}_{0}\text{-}prediction$ diffusion model [29], where the model outputs the denoised feature $\mathbf{x}_{0}$ based on the noisy feature $\mathbf{x}_{t_{i}}$, the time condition $t_{i}$ and the text condition $c$.
+According to the definition of Rectified Flow (RF) [19], there is the following transfer equation:
 $\mathbf{x}_{t_{i}}=t_{i}\mathbf{x}_{1}+(1-t_{i})\mathbf{x}_{0}.$ (15)
-According to the theory of stochastic interpolation (Albergo et al., 2023), RF effectively approximates $\mathbf{x}_{1}-\mathbf{x}_{0}$ by modeling $\mathbf{v}_{t_{i}}$:
+According to the theory of stochastic interpolation [1], RF effectively approximates $\mathbf{x}_{1}-\mathbf{x}_{0}$ by modeling $\mathbf{v}_{t_{i}}$:
 $\mathbf{v}_{t_{i}}=\mathbf{x}_{1}-\mathbf{x}_{0}.$ (16)
@@
 $\mathbf{x}_{\theta}(\mathbf{x}_{i},t_{i},c)=\mathbf{x}_{i}-\mathbf{v}_{\theta}(\mathbf{x}_{i},t_{i},c)\cdot t_{i}.$ (18)
-Taking the multistep second-order DPMSolver++ as an example (see Algorithm 2 in  (Lu et al., 2022b)), we derive the corrected $\mathbf{x}_{\theta}$ for the RF sampling process as $\mathbf{D}_{i}$:
+Taking the multistep second-order DPMSolver++ as an example (see Algorithm 2 in  [22]), we derive the corrected $\mathbf{x}_{\theta}$ for the RF sampling process as $\mathbf{D}_{i}$:
 $\mathbf{D}_{i}\leftarrow\left(1+\frac{h_{i}}{2h_{i-1}}\right)\left(\mathbf{x}_{i-1}-\mathbf{v}_{\theta}(\mathbf{x}_{i-1},t_{i-1},c)\cdot t_{i-1}\right)$ (19)
@@
 $\lambda_{t_{i}}:=\log\left(\frac{1-t_{i}}{t_{i}}\right).$ (20)
-Based on the exact discretization formula for the probability flow ODE proposed in DPM-Solver++ (Eq. (9) in (Lu et al., 2022b)), we can derive the final transfer equation:
+Based on the exact discretization formula for the probability flow ODE proposed in DPM-Solver++ (Eq. (9) in [22]), we can derive the final transfer equation:
 $\mathbf{x}_{i}\leftarrow\frac{t_{i}}{t_{i-1}}\mathbf{x}_{i-1}-(1-t_{i})\left(e^{-h_{i}}-1\right)\mathbf{D}_{i},\quad 1\leq i<T.$ (21)
-Appendix D MixGRPO Algorithm
-Taking the progressive movement strategy of the sliding window as an example, the MixGRPO Algorithm 1 is as follows.
-MixGRPO first sets the left boundary $l$ of the sliding window $W(l)$ at $t=0$ of the reverse process, which corresponds to the first sampling time-step.
-A GRPO iteration consists of two phases: sampling and optimization.
-In the sampling phase, mixed ODE-SDE sampling is performed for the $N$ samples within the group, using SDE within the window and ODE outside.
-Then, the reward and advantage for the $N$ samples are calculated.
-For the multi-reward strategy, the advantage for each reward model is computed separately and then aggregated to obtain the final advantage.
-In the optimization phase, the GRPO loss (Shao et al., 2024) is computed and optimized only at the sampling timesteps within the SDE window $W(l)$.
-After each GRPO iteration, the current training steps are checked against the conditions for moving the window, determining whether to shift the window by the window stride $s$.
-Algorithm 1 MixGRPO Training Process
-1:initial policy model $\pi_{\theta}$; reward models $\{R_{k}\}_{k=1}^{K}$; prompt dataset $\mathcal{C}$; total sampling steps $T$; number of samples per prompt NN;
-2:sliding window $W(l)$, window size $w$, shift interval $\tau$, window stride $s$
-3:Init left boundary of $W(l)$: $l\leftarrow 0$
-4:for training iteration $m=1$ to $M$ do
-5:  Sample batch prompts $\mathcal{C}_{b}\sim\mathcal{C}$
-6:  Update old policy model: $\pi_{\theta_{\text{old}}}\leftarrow\pi_{\theta}$
-7:  for each prompt $\mathbf{c}\in\mathcal{C}_{b}$ do
-8:   Init the same noise $\mathbf{x_{0}}\sim\mathcal{N}(0,\mathbf{I})$
-9:   for generate $i$-th image from $i=1$ to N do
-10:     for sampling timestep $t=0$ to T1T-1 do $\triangleright$ $\pi_{\theta_{\text{old}}}$ mixed sampling loop
-11:      if $t\in W(l)$ then
-12:        Use SDE Sampling to get $\mathbf{x}^{i}_{t+1}$
-13:      else
-14:        Use ODE Sampling to get $\mathbf{x}^{i}_{t+1}$
-15:      end if
-16:     end for
-17:   end for
-18:   for $i$-th image from $i=1$ to N do
-19:     Calculate multi-reward advantage: $A_{i}\leftarrow\sum_{k=1}^{K}\frac{R(\mathbf{x}^{i}_{T},\mathbf{c})^{i}_{k}-\mu_{k}}{\sigma_{k}}$
-20:   end for
-21:   for optimization timestep $t\in W(l)$ do $\triangleright$ optimize policy model $\pi_{\theta}$
-22:     Update policy model via gradient ascent: $\theta\leftarrow\theta+\eta\nabla_{\theta}\mathcal{J}$
-23:   end for
-24:  end for
-25:  if $m\bmod\tau~\text{is}~0$ then $\triangleright$ move sliding window
-26:   $l\leftarrow\min(l+s,~T-w)$
-27:  end if
-28:end for
-Appendix E MixGRPO-Flash Algorithm
+9 MixGRPO-Flash Algorithm
 MixGRPO-Flash Algorithm 2 accelerates the ODE sampling that does not contribute to the calculation of the policy ratio after the sliding window by using DPM-Solver++ in the Eq.
@@
 […]
-16:        Use higher-order ODE sampling to get $\mathbf{x}^{i}_{t+1}$
+16:         Use DPM-Solver++ sampling to get $\mathbf{x}^{i}_{t+1}$
 17:      end if
@@
 20:    for $i$-th image from $i=1$ to N do
-21:     Calculate multi-reward advantage: $A_{i}\leftarrow\sum_{k=1}^{K}\frac{R(\mathbf{x}^{i}_{\tilde{T}},\mathbf{c})^{i}_{k}-\mu_{k}}{\sigma_{k}}$
+21:     Calculate advantage: $A_{i}\leftarrow\sum_{k=1}^{K}\frac{R(\mathbf{x}^{i}_{\tilde{T}},\mathbf{c})^{i}_{k}-\mu_{k}}{\sigma_{k}}$
 22:    end for
-23:   for optimization timestep $t\in W(l)$ do $\triangleright$ optimize policy model $\pi_{\theta}$
-24:     Update policy model via gradient ascent: $\theta\leftarrow\theta+\eta\nabla_{\theta}\mathcal{J}$
+23:    for optimization timestep $t\in W(l)$ do
+24:     Update policy model: $\theta\leftarrow\theta+\eta\nabla_{\theta}\mathcal{J}$
 25:    end for
@@
 34:end for
-Appendix F More visualized results
-Figure 5: Qualitative comparison with different training-time sampling steps.
-The performance of MixGRPO does not significantly decrease with the reduction in overhead.
-The Frozen strategy means that optimization is only employed at the initial denoising steps.
+10 Hybrid Inference for Solving Reward Hacking
+As discussed in Section 5, reward hacking stems from the limited evaluation capabilities of the reward model.
+To address reward hacking and improve visualization, we employ the hybrid inference strategy from [43], which uses the post-trained model for low-SNR (signal-to-noise ratio) steps and the original model for high-SNR steps during inference-time sampling.
+In our experiments, we also applied hybrid inference to the other baseline models to ensure a fair and consistent comparison.
+We employ hybrid inference and introduce the hybrid percent $p_{\text{mix}}$.
+This means that the initial, high-SNR $p_{\text{mix}}T$ denoising steps, are handled by the model trained with GRPO, while the remaining denoising process is finished by the original model.
+Table 10 and Figure 5 respectively illustrate the changes in performance and images as $p_{\text{mix}}$ increases under the multi-rewards training scenario.
+The experimental results demonstrate that $p_{\text{mix}}=80\%$ is an optimal empirical value that effectively mitigates hacking while maximizing alignment with human preferences.
+Table 10: Comparison with different hybrid inference percentages
+$p_{\text{mix}}$ HPS-v2.1 Pick Score ImageReward Unified Reward
+0% 0.313 0.226 1.089 3.369
+20% 0.342 0.233 1.372 3.386
+40% 0.356 0.235 1.539 3.395
+60% 0.362 0.236 1.598 3.407
+80% 0.366 0.238 1.610 3.411
+100% 0.369 0.238 1.607 3.378
+Figure 5: Qualitative comparison with different hybrid inference percentages. We found 80% to be the optimal value, as it maximizes image quality without causing over-optimization.
+11 Cross-Dataset Experiments
+To investigate the robustness and parameter sensitivity of the sliding window strategy in MixGRPO, we conducted a series of cross-dataset ablation studies.
+We established two reciprocal settings to evaluate both in-domain (ID) and out-of-domain (OOD) performance.
+In cross-dataset experiment 1, the model was trained on the HPDv2 [40] dataset and evaluated on the test sets of HPDv2 (ID) and Pick-a-Pic v1 (OOD).
+In cross-dataset experiment 1, conversely, the model was trained on Pick-a-Pic v1 [12] and evaluated on the test sets of Pick-a-Pic v1 (ID) and HPDv2 (OOD).
+Within these settings, we ablated the key parameters of MixGRPO: the moving strategy, shift interval $\tau$, window size $w$ and window stride $s$.
+The results are presented in the Tables 11121314.
+As shown in Table 11, Progressive-Decay(Exp) and Progressive-Constant emerge as the top-performing moving strategies.
+While Progressive-Decay(Exp) consistently achieves the highest ImageReward [41] score in in-domain evaluations, Progressive-Constant demonstrates superior overall performance.
+Notably, the Progressive-Constant strategy consistently outperforms all others in out-of-domain tests, demonstrating its strong robustness and generalization capabilities.
+The results concerning the shift interval ($\tau$), which dictates the frequency of window movement in training iteration steps, are summarized in Table 12.
+Overall, the optimal performance is consistently observed at $\tau=25$.
+Specifically, as $\tau$ is incrementally increased from $15$ to $25$, performance metrics exhibit a gradual rise across both in-domain and out-of-domain test sets.
+However, a crucial observation is the sharp decline in performance when the interval is extended to $\tau=30$, with results even falling below the baseline achieved at $\tau=15$.
+This trend suggests that a moderate reduction in the sliding window’s movement speed (from $\tau=15$ to $\tau=25$) allows the temporal behavior within the window to be sufficiently optimized.
+Conversely, overly slow window movement (exceeding $\tau=25$) is prone to over-optimization at certain timesteps, causing the model’s distribution to diverge significantly from the target preferences of the reward models.
+The cross-dataset experiments further confirm the universality of $\tau=25$ as the optimal setting.
+The comparative results for the window size ($w$) are presented in Table 13.
+It is evident that both $w=4$ and $w=6$ represent optimal settings.
+Specifically, $w=6$ slightly surpasses $w=4$ only on the in-domain test set of Cross-dataset Experiment 1, and in other experiments, it only holds a marginal advantage in the HPS-v2.1 [40] score.
+It is crucial to note that the window size $w$ corresponds to the number of denoising timesteps that must be optimized during each training iteration, which, in turn, is linearly correlated with the overall optimization overhead of the Reinforcement Learning (RL) process.
+Consequently, $w=4$ is identified as the best compromise, offering an optimal trade-off between performance and computational efficiency.
+Regarding parameter sensitivity, we observe that the change in performance metrics—across both in-domain and out-of-domain tests—is relatively minimal when $w$ varies between $2$ and $6$.
+This stability suggests that the model exhibits robust performance with respect to the window size setting.
+As long as the window size remains within a reasonable range (e.g., $2\leq w\leq 6$), the model is capable of effective learning and maintaining stable performance, thereby reducing necessity for meticulous hyperparameter tuning.
+The results concerning the stride ($s$), which defines the step size for each sliding window movement, are presented in Table 14.
+Considering all experimental metrics across the different datasets, a stride of $s=1$ is identified as the optimal setting overall.
+It is important to note that the optimal window size of $w=4$ was used in this experiment.
+With $s=1$, the last three timesteps within the window are repeatedly optimized in the subsequent training cycle.
+Crucially, the experimental results indicate that this repetition leads to performance improvement.
+This is likely due to the nature of the high-SNR denoising process, where the initial timesteps (e.g., the first timestep in the window) are subject to a greater extent of GRPO optimization, despite the window size being set at $w=4$.
+Conversely, the behavior learned during the subsequent three timesteps remains relatively under-optimized (underfitted) and thus significantly benefits from this repeated optimization.
+Furthermore, we observe that the highest in-domain HPS-v2.1 [40] score is achieved when $s=3$.
+This localized optimal result may be attributed to the HPS-v2.1 [40] reward model’s preference alignment being relatively easier to optimize, suggesting that repeated optimization (i.e., smaller strides) could potentially induce over-optimization in this specific metric.In summary, $s=1$ proves to be the most robust and generalizable choice across the varying experimental conditions.
+Table 11: Comparison for different moving strategies in cross-dataset experiments. The optimal setting is highlighted in green.
+Strategy Interval Schedule Cross-dataset Experiment 1 Cross-dataset Experiment 2
+In Domain Dataset (HPD v2) Out-of-Domain Dataset (Pick-a-Pi v1) In Domain Dataset (Pick-a-Pi v1) Out-of-Domain Dataset (HPD v2)
+HPS-v2.1 Pick Score ImageReward HPS-v2.1 Pick Score ImageReward HPS-v2.1 Pick Score ImageReward HPS-v2.1 Pick Score ImageReward
+Frozen / 0.354 0.234 1.580 0.352 0.226 1.539 0.346 0.230 1.601 0.351 0.231 1.587
+Random Constant 0.365 0.237 1.513 0.361 0.230 1.512 0.349 0.227 1.524 0.343 0.225 1.530
+Progressive Decay(Linear) 0.365 0.235 1.566 0.363 0.229 1.572 0.363 0.231 1.614 0.355 0.230 1.597
+Decay(Exp) 0.360 0.239 1.632 0.364 0.232 1.612 0.361 0.234 1.628 0.358 0.230 1.584
+Constant 0.367 0.237 1.629 0.366 0.234 1.622 0.365 0.238 1.618 0.359 0.232 1.601
+Table 12: Comparison for different shift intervals $\tau$ in cross-dataset experiments. The optimal setting is highlighted in green.
+$\tau$ Cross-dataset Experiment 1 Cross-dataset Experiment 2
+In Domain Dataset (HPD v2) Out-of-Domain Dataset (Pick-a-Pi v1) In Domain Dataset (Pick-a-Pi v1) Out-of-Domain Dataset (HPD v2)
+HPS-v2.1 Pick Score ImageReward HPS-v2.1 Pick Score ImageReward HPS-v2.1 Pick Score ImageReward HPS-v2.1 Pick Score ImageReward
+15 0.366 0.237 1.509 0.359 0.227 1.470 0.358 0.234 1.610 0.353 0.228 1.542
+20 0.366 0.238 1.610 0.360 0.228 1.619 0.361 0.237 1.615 0.357 0.233 1.588
+25 0.367 0.237 1.629 0.366 0.234 1.623 0.365 0.238 1.618 0.359 0.232 1.601
+30 0.350 0.229 1.589 0.348 0.221 1.585 0.355 0.234 1.609 0.351 0.229 1.509
+Table 13: Comparison for different window sizes $w$ in cross-dataset experiments. The optimal setting is highlighted in green.
+$w$ Cross-dataset Experiment 1 Cross-dataset Experiment 2
+In Domain Dataset (HPD v2) Out-of-Domain Dataset (Pick-a-Pi v1) In Domain Dataset (Pick-a-Pi v1) Out-of-Domain Dataset (HPD v2)
+HPS-v2.1 Pick Score ImageReward HPS-v2.1 Pick Score ImageReward HPS-v2.1 Pick Score ImageReward HPS-v2.1 Pick Score ImageReward
+1 0.359 0.232 1.571 0.349 0.223 1.445 0.349 0.229 1.553 0.346 0.221 1.565
+2 0.366 0.235 1.618 0.363 0.230 1.614 0.363 0.234 1.597 0.359 0.230 1.585
+4 0.367 0.237 1.629 0.366 0.234 1.623 0.365 0.238 1.618 0.359 0.232 1.601
+6 0.370 0.238 1.624 0.364 0.232 1.620 0.366 0.237 1.608 0.362 0.232 1.594
+Table 14: Comparison for different window strides $s$ in cross-dataset experiments. The optimal setting is highlighted in green.
+$s$ Cross-dataset Experiment 1 Cross-dataset Experiment 2
+In Domain Dataset (HPD v2) Out-of-Domain Dataset (Pick-a-Pi v1) In Domain Dataset (Pick-a-Pi v1) Out-of-Domain Dataset (HPD v2)
+HPS-v2.1 Pick Score ImageReward HPS-v2.1 Pick Score ImageReward HPS-v2.1 Pick Score ImageReward HPS-v2.1 Pick Score ImageReward
+1 0.367 0.237 1.629 0.366 0.234 1.623 0.364 0.238 1.618 0.359 0.232 1.601
+2 0.357 0.236 1.575 0.357 0.233 1.587 0.363 0.234 1.574 0.354 0.231 1.591
+3 0.370 0.236 1.578 0.364 0.230 1.579 0.368 0.237 1.586 0.358 0.228 1.585
+4 0.368 0.238 1.575 0.349 0.224 1.573 0.366 0.231 1.566 0.357 0.225 1.568
+12 More Visualized Results
 Figure 6: Comparison of the visualization results of FLUX, DanceGRPO, and MixGRPO under HPS-v2.1 as the reward model.

### v4

@@
 Specifically, we employ a mixed ODE-SDE strategy, applying SDE sampling to a denoising sub-interval and Ordinary Differential Equations (ODE) sampling to the rest.
+Recently, Coefficients-Preserving Sampling (CPS) [37] was proposed as a more principled alternative to standard SDE sampling.
+Further details are provided in 12.
 Figure 2: (Left) Methods Comparison.
@@
 According to Eq.
-(1), we only need to optimize GRPO [32] at the time within the interval $S$ for $N$ samples in the group.
+(1), we only need to optimize GRPO [32] at the time within the interval $S$ for NN samples in the group.
 For an unbiased advantage estimation in GRPO, we constrain the source of variation across images generated from a single prompt.
@@
 Mixed precision training is used with the bfloat16 (bf16) format, while the master weights are maintained at fp32 precision.
+We also tried CPS sampling instead of normal SDE sampling, the details can seen in 12 .
 4.3 Comparison with existing methods
@@
 Consequently, $w=4$ is identified as the best compromise, offering an optimal trade-off between performance and computational efficiency.
-Regarding parameter sensitivity, we observe that the change in performance metrics—across both in-domain and out-of-domain tests—is relatively minimal when $w$ varies between $2$ and $6$.
+Regarding parameter sensitivity, we observe that the change in performance metrics—across both in-domain and out-of-domain tests—is relatively minimal when $w$ varies between 22 and 66.
 This stability suggests that the model exhibits robust performance with respect to the window size setting.
@@
 4 0.368 0.238 1.575 0.349 0.224 1.573 0.366 0.231 1.566 0.357 0.225 1.568
+12 Coefficients-Preserving Sampling
+In our MixGRPO framework, introducing stochasticity during the inference phase is crucial for effective exploration in reinforcement learning.
+While a common practice involves the use of Stochastic Differential Equations (SDEs), we adopt Coefficients-Preserving Sampling (CPS) [37] as a more refined alternative to maintain the integrity of the probability path.
+The standard SDE-based sampling, often discretized through the Euler-Maruyama method, follows the update rule:
+$\mathbf{x}_{t_{i-1}}=\mathbf{x}_{t_{i}}-\mathbf{v}_{t_{i}}(\mathbf{x}_{t_{i}},t_{i})\Delta t+\sqrt{2\sigma^{2}\Delta t}\mathbf{\epsilon}_{i},\qquad\mathbf{\epsilon}_{i}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ (24)
+Although this formulation provides the necessary stochasticity, it tends to inject excessive independent noise at each discretization step, resulting in ”grainy” artifacts in the generated samples.
+As observed in our preliminary experiments, such high-frequency noise can lead to reward hacking, where reward models (e.g., Pick Score [12] or HPS-v2.1 [41]) prioritize low-level textures over structural coherence, thereby hindering the convergence of RL optimization.
+To address these issues, we utilize CPS [37], which reformulates the transition process by drawing inspiration from DDIM [33] framework. The update rule for CPS [37] is defined as:
+$\mathbf{x}_{t_{i-1}}=\frac{1-t_{i-1}}{1-t_{i}}\mathbf{x}_{t_{i}}+\left(t_{i-1}-\frac{1-t_{i-1}}{1-t_{i}}t_{i}\right)\mathbf{v}_{t_{i}}+\sigma_{t_{i}}\mathbf{\epsilon}_{i}$ (25)
+The core advantage of CPS [37] lies in its strategic construction of coefficients for $\mathbf{x}_{t_{i}}$ and $\mathbf{v}_{t_{i}}$, which preserves the linear interpolation structure of Flow Matching while introducing controlled stochasticity via $\sigma_{t_{i}}$.
+By effectively eliminating sampling artifacts inherent in SDE-based methods, CPS [37] yields cleaner images that provide more reliable feedback for MixGRPO.
+In our implementation, we set $NFE_{\pi_{\theta_{old}}}=25$, window size $w=4$, and window stride $s=1$ with fixed initial noise, employing HPS-v2.1 [41], Pick Score [12], and ImageReward [42] as multi-reward metrics.
+Both the original SDE sampling and CPS [37] sampling were trained for 300 steps and evaluated on the HPDv2 dataset [41].
+As demonstrated in Table 15, the integration of CPS [37] leads to superior performance across all reward metrics, with qualitative comparisons in Figure 9 further illustrating its enhanced visual fidelity.
+Table 15: Comparison with CPS sampling
+Method Pick Score ImageReward HPS-v2.1
+FLUX 0.2265 1.0883 0.3126
+MixGRPO-SDE 0.2344 1.5901 0.3648
+MixGRPO-CPS 0.2382 1.6449 0.3692
 13 More Visualized Results
@@
 Figure 8: Comparison of the visualization results of SD3.5-M, offline DPO, online DPO, Flow-GRPO and MixGRPO under HPS-v2.1, Pick Score and ImageReward as multi-reward models.
+Figure 9: Comparison of the visualization results of FLUX, SDE sampling and CPS sampling